### v1

# Spin-glass dynamics: experiment, theory and simulation

Abstract
The study of spin-glass dynamics, long considered the paradigmatic complex system, has reached important milestones. The availability of high-quality single crystals has allowed the experimental measurement of spin-glass coherence lengths of almost macroscopic dimensions, while the advent of special-purpose massive computers (the Janus Collaboration) enables dynamical simulations that approach experimental time and length scales. This review provides an account of the quantitative convergence of these two avenues of research, with precise experimental measurements of the expected scaling laws and numerical reproduction of classic experimental results, such as memory and rejuvenation. The article opens with a brief review of the defining spin-glass properties, randomness and frustration, and their experimental consequences. These apparently simple characteristics are shown to generate rich and complex physics. Models are introduced that enable quantitative dynamical descriptions, either analytically or through simulations. The many theoretical pictures of the low-temperature phase are reviewed. After a summary of the main numerical results in equilibrium, paying particular attention to the concept of temperature chaos, this review examines off-equilibrium dynamics in the absence of a magnetic field and shows how it can be related to the structure of the equilibrium spin-glass phase through the fluctuation-dissipation relations. The nonlinear response at a given temperature is then developed, including experiments and scaling in the vicinity of the spin-glass transition temperature $T_{\mathrm{g}}$. The consequences of temperature change —including temperature chaos, rejuvenation, and memory— are reviewed. The interpretation of these phenomena requires the identification of several length scales relevant to dynamics, which, in turn, generate new insights. Finally, issues for future investigations are introduced, including what is to be “nailed down” theoretically, why the Ising Edwards-Anderson model is so successful at modeling spin-glass dynamics, and experiments yet to be undertaken.
I Introduction (and a little history)
The study and broad applications of spin-glass dynamics have advanced remarkably since the last major review [63], nearly forty years ago. This work aims to bring the experimental and theoretical accomplishments up to date and point to future opportunities. While many individual summaries are available [218, 296, 78, 83, 9, 281], this review combines experimental and theoretical perspectives in a synergistic format. The give and take between experiment, theory, and simulations caused by these interactions is a perfect example of the “three legs of scientific research.” The accomplishments of the Uppsala and Saclay groups were responsible for the surge of interest in the field in the late 20th and early 21st century. The following advent of powerful computational facilities custom designed for Ising spin glasses produced simulations, both in and out of equilibrium, that have enabled direct comparisons with experimental results. These in turn have inspired new experiments that have, in combination with simulations, led to a deep quantitative understanding of spin-glass dynamics.
I.1 What is a spin glass?
Over the years since the original experiments of Cannella and Mydosh [74], the combination of randomness mixed with frustration [274] has been identified as the essential property of a spin glass [15]. Simply defined, frustration means that not all exchange-coupled spin orientations can be energetically satisfied for any configuration. While each property by itself results in a different class (e.g., an antiferromagnet in a triangular lattice is fully frustrated), the combination of the two is defining for spin glasses so far as we know. Spin-glass behavior is found in both metals and insulators with competing ferro- and anti-ferromagnetic couplings. The RKKY oscillating coupling [255, 165, 295] in dilute magnetic alloys [e.g., CuMn, 132, 164] and competing exchange couplings in insulators [e.g., CdCr1.7In0.3S4, 281] are stereotypical examples.
What is so wonderful about this “simple” set of requirements is that [16]: “the key result here is the beautiful revelation of the structure of the randomly ‘rugged landscape’ that underlies many complex optimization problems […] Physical spin glasses and the SK model [260] are only a jumping-off point for an amazing cornucopia of wide-ranging applications of the same kind of thinking.”
The purpose of this review is to explore the dynamics of the spin-glass state. The richness of the ensuing consequences will prove remarkable. What will be even more impressive is the applicability of these results to a great variety of physical and even sociological systems. As noted by Parisi [240], there is an interplay between disorder and fluctuations in physical systems from atomic to planetary scales. Theories based on multiple equilibria are present in many disciplines: the “punctuated equilibria” of evolutionary theory [105]; in ecosystems, glaciations and geological eras; memory through associative neural networks [150]; the physics of structural glasses [131]; the multiple equilibria of economics [103], etc.
The utility of solutions of the spin-glass problem is precisely this. Over the past five decades, from their experimental properties, together with somewhat simplified theoretical models (though with highly complex mathematics!), the field has amassed Anderson’s “cornucopia” of concepts, solutions, and insights. The authors of this review hope that, by exhibiting these results, the reader will be able to understand and attack other complex systems of interest over a wide variety of applications.
I.2 Experimental milestones
Interest in the magnetic alloys now known as spin glasses began with the work of Cannella and Mydosh, which showed a cusp in the ac susceptibility of dilute Au:Fe alloys [74, 75, see Fig. 1]. The sharpness of the cusp indicated what might be a phase transition, while earlier dc magnetization measurements on similar alloys only showed a rounded transition [see, for example, 195].
Figure 1: Susceptibility for AuFe $1\leq{\text{c}}\leq 8$ at.in the region of low magnetic fields. The data were taken every $0.25$ K around the peak and every $0.5$ or 1 K elsewhere. The scatter of the points is of the order of the thickness of the lines. The open circles indicate isolated points taken at higher temperatures. Reproduced from Fig. 9 of Cannella and Mydosh [74].
A simple picture described the system as merely a distribution of independent magnetic domains, so that the cusp was an artifact, caused by the freezing of a portion of the sample at a given temperature and frequency [271, 272, 119]. In such a case, a variation in the measurement frequency would significantly change the temperature of the cusp. This question was addressed by the UCLA group [92, 91]. They measured the frequency dependence of the cusp temperature and found the maximum shift to be of the order of 50 mK from 16 Hz to 2.8 MHz (Fig. 2). This was strong evidence that the cusp was the signature of some type of phase transition.
[…]
In the early years, prior to this study in an extended frequency range, the dc magnetic measurements had not revealed critical behavior. The search was on for another signature, and it seemed obvious that heat capacity should reveal some sort of thermodynamic signature. However, heat-capacity measurements by Wenger and Keesom [289, 290] showed no evidence (see Fig. 3).
Figure 3: Specific heat of Au0.92Fe0.08 in the temperature region 3-50 K is shown as a plot of $C/T$ vs $T^{2}$. The solid curve is the calculated non-magnetic contribution to the specific heat of the alloy between 0 and 30 K. The inset shows the susceptibility results of the same sample, which were provided by S. A. Werner. Reproduced from Fig. 2 of Wenger and Keesom [289] and Wenger and Keesom [290].
Other measurements were made in hopes of observing something indicating critical behavior. They included neutron scattering, resistivity, Hall, thermopower, ultrasound, muon spin resonance, ultrasound, and Mossbauer measurements, but none provided the same strong signature as ac susceptibility [see 119, for an in depth discussion of each of these probes]. In 1991, it became clear that a thermodynamic phase transition took place at $T_{\mathrm{g}}$ [141]. This work was followed by simulations [233, 38] that supported the conclusion that spin glasses experience a continuous phase transition.
The spin-glass transition temperature has been defined as the point at which there is an onset of irreversibility, which is seen by comparing two measurements of the spin-glass magnetization: lowering the temperature in the presence of a magnetic field [field-cooled magnetization (FC)]; or lowering the temperature in the absence of a magnetic field and then turning it on [zero-field-cooled magnetization (ZFC)]. Fig. 4 displays the difference, the so-called thermoremanent magnetization (TRM), between both protocols [102]. Experimentally, the FC magnetization has a weak time dependence, while ZFC and TRM measurements are strongly time dependent.
It is tempting to conclude, because of the weak time dependence of the FC magnetization, that the sum of the time-dependent ZFC + TRM magnetizations should always equal the FC magnetization. This led the Uppsala group to pose the “extended principle of superposition” [228, 193, 227, 100],
$M_{{\text{TRM}}}(t,t_{\mathrm{w}})+M_{{\text{ZFC}}}(t,t_{\mathrm{w}})=M_{{\text{FC}}}(0,t_{\mathrm{w}}+t),$ (1)
where $t_{\mathrm{w}}$ is the waiting or aging time (see glossary in Sec. I.3). However, Paga et al. [231] showed that Eq. (1) is valid only for very small $H$ (see Sec. V.2).
Among the numerous time-dependent studies, there are some key elements that deserve mentioning. The first is the quasi-logarithmic time dependence of the magnetization, first observed in AuFe alloys by Tournier [275], see Fig. 5, and in FeCr by Ishikawa et al. [154].
[…]
Figure 5: Time decay of the thermoremanent magnetization (TRM) of Au92Fe8. Reproduced from Tournier [275].
Figure 6: Field dependence of the thermoremanent magnetization (TRM) of AuFe 0.5 at.obtained after cooling from $T>T_{\mathrm{g}}$ to $T=1.2$ K in a magnetic field $H$ which is them removed and of the isothermal remanent magnetization (IRM) obtained when cooling in zero magnetic field, turning on the magnetic field $H$ at 1.2 K and then removing it. Reproduced from Fig. 4 of Tholence and Tournier [271].
Early on, the TRM and the isothermal remnant magnetization (IRM) were most commonly measured [for example, 271, see Fig. 6]. An IRM measurement consists of cooling to a temperature below the glass temperature in zero applied magnetic field. A magnetic field is applied and removed with the TRM being the remaining magnetization.
A useful tool for quantitative extraction of spin-glass dynamics has been the measurement of $S(t,t_{\mathrm{w}};H)$,
$S(t,t_{\mathrm{w}};H)=(\pm){\frac{\text{d}M(t_{\mathrm{w}},T;H)}{\text{d}\ln t}}\,,$ (2)
where $+$ (--) refers to ZFC (TRM) measurements. This was first developed by Nordblad et al. [228], who showed that the time for $S(t,t_{\mathrm{w}};H)$ to peak was a useful measure of a spin-glass response time, $t_{\mathrm{w}}^{\text{eff}}$ (see Fig. 7). Typically, $t_{\mathrm{w}}^{\text{eff}}\approx t_{\mathrm{w}}$, but there are important exceptions [174] for temperatures close to $T_{\mathrm{g}}$.
Figure 7: Relaxation rates $S(t)$ of the remanent magnetization of CuMn 5 at.($T_{\mathrm{g}}=28$ K) at $T=21$ K for different waiting times $t_{\mathrm{w}}$. A relaxation rate of 0.1% of the field-cooled susceptibility value is indicated on the figure. Reproduced from Fig. 2 (c) of Nordblad et al. [228].
Figure 8: The correlation length $\xi$ against the age $t_{\text{w}}$ (log scale) at different temperatures $T_{\mathrm{m}}-\Delta T,T_{\mathrm{m}},$ and $T_{\mathrm{m}}+\Delta T$. The sample was CuMn 10 at.%, with $T_{\mathrm{g}}=45.2$ K, and $T_{\mathrm{m}}=41.2$ K. Granberg et al. [134] claimed that the different wait times $t_{\text{w}1},t_{\text{w}2}$, and $t_{\text{w}3}$ yielded the same value $\xi_{0}$ of the correlation length. Reproduced from Fig. 6 of Granberg et al. [134].
Later work by the Uppsala group [134] used temperature shifts to explore the $S(t)$ peaks as a function of wait times to determine how the growth rate of the correlation length depends upon temperature. They assumed the correlation length scaled with the time $t$ and argued that at a given temperature, the correlation length (their $\xi_{0}$) was wait-time independent. Their data is summarized in Fig. 8.
Figure 9: A plot of $N_{\text{s}}$, the number of spins participating in barrier quenching (and barrier hopping), for CuMn 6 at.vs $t_{\mathrm{w}}$ on a log scale at fixed $T=0.89\,T_{\mathrm{g}}=28$ K. The solid curve drawn through the points is the prediction for power-law dynamics [182], while the dashed curve is the prediction for activated dynamics [115, 114] with the exchange factor set equal to $T_{\mathrm{g}}$ (i.e., independent of $T$ and $t$). The two fits are equally good. Reproduced from Fig. 3 of Joh et al. [159].
Following this work, Joh et al. [159] used the field dependence of the $S(t,t_{\mathrm{w}};H)$ peaks as a function of waiting times to quantify a correlation length of the spin-glass state, $\xi(t_{\mathrm{w}},T)$. They were able to extract an explicit value for $\xi(t_{\mathrm{w}},T)$ —the $\xi_{0}$ of Fig. 8— and its growth rate as a function of $t$, $t_{\mathrm{w}}$ and $T$. Setting the number of correlated spins $N_{\text{s}}\approx\xi(t_{\mathrm{w}},T)^{3}$, they measured $\xi(t_{\mathrm{w}},T)$ as a function of $t_{\mathrm{w}}$. Their results for CuMn 6 at.are exhibited in Fig. 9. Note the different dependence upon $t_{\mathrm{w}}$ between $\xi_{0}$ and $\xi(t_{\mathrm{w}},T)^{3}$ in Figs. 8 and 9, respectively.
The development of the ability to determine an experimental value for the correlation length was critical for connecting experiments to models and simulations. The rationale lies in the dynamics of a spin glass. The defining physical property of glassy materials is determined by the energy landscape and that, in turn, is created by the growth of the correlated domains.
Progress in the experimental state has largely been based on three physical properties intrinsic to the spin-glass state. They are aging, rejuvenation, and memory. The aging phenomenon has been mentioned earlier beginning with the work of Tournier [275]. The discovery of rejuvenation and memory was first reported by Jonason et al. [162] using low-frequency ac susceptibility measurements. This experiment will be discussed in detail in Sec. VI.2 as the basis for more recent work on memory and rejuvenation.
Figure 10: Barrier distributions, $D(\Delta)$, at $T=0$, based on measurements of $S_{R}(f,T)$ for five film thicknesses: (a) 10 nm, (b) 18 nm, (c) 25 nm, (d) 40 nm, and (e) 80 nm. The bold portions fall within our measurement bandwidth. The dashed portions are an extrapolation to guide the eye. Reproduced from Fig. 3 of Harrison et al. [144].
Using the development of an experimental quantification of the memory effect, there has been recent progress in understanding its origin in terms of temperature-specific correlations [125, 230].
Given that the energy barriers, $\Delta$, are critical to the dynamics, there have been attempts to determine their distribution, $D(\Delta)$, for the spin-glass state. Recent $1/f$ electronic-noise measurements have been found to provide a measure of the energy-barrier distribution in thin films [144, 276]. Although a distribution was obtained, the relevant length scale was that of the electron mean free path and thus may be slanted towards the larger barriers (Fig. 10).
I.3 A glossary for aging dynamics
Important concepts used repeatedly in this review:
  1. 1.
    Two relevant temperatures are $T_{\mathrm{g}}$ and $T_{\mathrm{m}}$. The glass transition occurs at $T_{\mathrm{g}}$ (which also is the critical temperature for the phase transition that separates the high-temperature paramagnetic phase from the spin-glass phase that becomes stable below $T_{\mathrm{g}}$). In a typical experimental setting, the system is quickly cooled from some very high temperature $T>T_{\mathrm{g}}$ to the measuring temperature $T_{\mathrm{m}}$.
  2. 2.
    Two relevant experimental protocols, explained in Sec. I.2, are named zero-field cooling (ZFC) and thermoremanent magnetization (TRM).
  3. 3.
    Two relevant times are $t$ and $t_{\mathrm{w}}$. We set our initial time as the time for the temperature quench from $T>T_{\mathrm{g}}$ to $T_{\mathrm{m}}<T_{\mathrm{g}}$. The system is allowed to relax at $T_{\mathrm{m}}$ for the waiting time $t_{\mathrm{w}}$. Then, at time $t_{\mathrm{w}}$ the value of the externally applied magnetic field is varied (in the ZFC protocol the field is switched on at time $t_{\mathrm{w}}$; instead, in the TRM protocol the field is switched off). The magnetization is measured at a later time $t+t_{\mathrm{w}}$, giving us information about the sample’s reaction to the variation of the magnetic field at the earlier time $t_{\mathrm{w}}$.
  4. 4.
    We shall need two kind of averages. The thermal average of a quantity $\mathcal{A}$ will be denoted by $\langle\mathcal{A}\rangle$. The average over disorder, which is always performed after the thermal average, will be denoted by $\overline{\langle\mathcal{A}\rangle}$. Although the average over disorder is absolutely crucial for numerical simulations, experiments are usually carried out for just one sample. Due to self-averaging, this is usually a sound procedure under experimental conditions (see Sec. II.7).
  5. 5.
    The coherence length $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ is the size of the slowly growing glassy domains. See Sec. IV.2 for details [e.g., Eq. (23)]. If the growth of $\xi_{\mathrm{micro}}$ is not isotropic, such as in a film geometry, we consider the perpendicular direction, $\xi_{\mathrm{micro}}^{\perp}$, and the parallel direction, $\xi_{\mathrm{micro}}^{\parallel}$ (see Sec. V.2).
  6. 6.
    The two-time correlation length $\zeta(t,t_{\mathrm{w}})$ describes the growth of order in the aging regime when the two times $t_{\mathrm{w}}$ and $t+t_{\mathrm{w}}$ are compared [157, 20, 44]. This quantity, defined in Sec. IV.2, will be vital for understanding temperature-varying protocols.
  7. 7.
    The experimentally accessible $\xi_{\mathrm{Zeeman}}$ introduced by Joh et al. [159] will be explained in Sec. 13. This length is related to the lowering of free-energy barriers because of the Zeeman effect and is non-linear in the magnetic field.
As explained in Sec. V, the relationship between the three length scales $\xi_{\mathrm{micro}}$, $\xi_{\mathrm{Zeeman}}$ and $\zeta$ is the key to the study of the non-linear response of a spin glass in a magnetic field. Although $\xi_{\mathrm{micro}}$ and $\xi_{\mathrm{Zeeman}}$ turn out to behave very similarly whenever they can be compared (which could suggest also referring to $\xi_{\mathrm{Zeeman}}$ as a coherence length), we shall follow the experimental tradition of referring to $\xi_{\mathrm{Zeeman}}$ as a correlation length.
II Preliminaries
Before getting into detailed analyses of spin-glass dynamics, it is useful to establish the “ground rules.” These are the basic tools, both theoretical and experimental, that are used to elucidate the “amazing cornucopia” anticipated by Anderson [16].
II.1 The Edwards-Anderson and Sherrington-Kirkpatrick models and overlap as an order parameter
The large impact of the Cannella and Mydosh [74] experimental results, describing a phase transition with novel features, soon made it a relevant goal to establish a theoretical description and, if possible, a complete understanding of the behavior of diluted solutions of, say, Mn in Cu, which just started to be called spin glasses.
By assuming that $\sum_{i,j}J_{ij}\epsilon_{ij}=0$ on all length scales of the system (where $J_{ij}$ is the strength of the quenched interaction between sites $i$ and $j$ and $\epsilon_{ij}$ is the probability of finding two interacting spins at those positions), Edwards and Anderson [104] introduced a very simplified model of classical spins (which will turn out to embed all relevant features of the problem), the Edwards-Anderson (EA) model. A different approach, based on the presence of potentially magnetized domains, had been proposed before by Adkins and Rivier [2]. The fact that the disordered interactions are distributed at random and cannot change in time is a crucial element of the model.
The EA model provided an Ising-model-like description of spin glasses. The Hamiltonian is
$\mathcal{H}_{\text{\tiny EA}}\equiv-\sum_{i,j}\phantom{}^{{}^{\prime}}\;J_{ij}\sigma_{i}\sigma_{j}+H\sum_{i}\sigma_{i}\;,$ (3)
where the primed sum runs over the pairs of nearest neighbors in a $D$-dimensional hypercubic lattice. The spins $\sigma_{i}$ can take different forms. For instance, in the Heisenberg spin glass, they are unit vectors in space. Unless otherwise specified, however, in this review, the EA model will always be considered with Ising spins that can only take the values ±$\pm 1$. The applied magnetic field $H$ will be considered to be zero, for now.
Notice that this Hamiltonian is written in natural units where $k_{\text{B}}=1$. To compare with experimental results, temperatures can be expressed in terms of $T_{\mathrm{g}}$ (more sophisticated scaling variables will be introduced later). As for the magnetic field, it can be estimated [33, 21], that $H=1$ roughly corresponds to $5\times 10^{4}$ Oe. Finally, the natural unit of time in simulations is the lattice sweep, which is equivalent to $\sim 1$ ps [218].
It is crucial that the quenched couplings, which do not change in the course of the dynamics, can have both signs. A possible choice is $J_{ij}=\pm 1$ with a probability of one half, or that they be normally distributed, with zero mean and unitary variance, $P\left(J_{ij}\right)\sim\exp(-{J_{ij}^{2}}/{2J^{2}})$. We will not discuss here the possibility of a non-zero expectation value for the couplings.
Because the quenched couplings can be both positive or negative and are distributed at random, some of the products of the couplings on a closed path of consecutive lattice links will be positive and some will be negative. This implies, in turn, that not all interactions can be satisfied at the same time; the system is frustrated, in the sense that all interactions can never be completely satisfied. The energy cannot be as low as it could be in a uniform system. Typically frustration leads to degeneracy: the low-energy states of such a system are not as good as those of a pure systems, but are numerous.
In the EA model, because the interactions between spins oscillate in sign as a function of separation, ferromagnetism does not appear at any relevant length scale. On the contrary, for $T$ lower than a critical $T_{\mathrm{g}}$, a spin-glass phase appears, with the spins staying close to fixed directions that appear to be random. This results in overall zero magnetization, but a non-zero square of the local magnetization and leads to the cusp in the susceptibility that was observed in experiments (see Sec. I.2).
Figure 11: Illustration of the spin overlap. The left panels show configurations of two real replicas of the same spin glass, which have evolved for a long time in the broken-symmetry phase in a Monte Carlo simulation. The color map shows the average magnetization in the XY plane, averaged over the Z axis, which is always zero. In contrast, the right panel shows the spin overlap of Eq. (5), where a clear pattern is visible. Figure taken from Baity-Jesi et al. [35].
The absence of long-range magnetic order leads to the concept of spin overlap as an order parameter, which will be crucial for all further development of the field. One considers the spin $\sigma_{i}$ at site $i$ and defines:
$q_{i}\left(t,t+\tau\right)\equiv\lim_{\tau\to\infty}\left\langle\sigma_{i}\left(t\right)\sigma_{i}\left(t+\tau\right)\right\rangle,$ (4)
[…]
$q=\frac{1}{N}\sum_{i=1}^{N}\sigma^{(1)}_{i}\sigma^{(2)}_{i}.$ (5)
This is the most common observable in numerical simulations. The concept of a non-zero overlap of configurations with zero magnetization is illustrated in Fig. 11.
After averaging $q$ over the disorder realizations, we can also introduce its probability density function, denoted by $P(q)$ (see Sec. III.1 for some examples). In addition, we can define the link overlap, which can be written as
$q_{\text{l}}=\frac{1}{ND}\sum\phantom{}^{{}^{\prime}}\;\sigma^{(1)}_{i}\sigma^{(2)}_{i}\sigma^{(1)}_{j}\sigma^{(2)}_{j}\,.$ (6)
As with the overlap, we can associate a $P(q_{\text{l}})$ to this observable. The behavior of the overlap and the link overlap under spin flip is very different: consider two configurations of spins with the same disorder which differ in a domain of size $L$, where the spins have flipped, the link overlap provides an estimate for the surface of the domain, while the overlap will reflect its volume.
The EA spin glass, an Ising-like finite-dimensional model with quenched, frustrated interaction, cannot be solved exactly and cannot be dissected too deeply by analytical techniques. The experience of the Ising ferromagnet suggested introducing the mean-field, non-local Sherrington-Kirkpatrick (SK) model [260, 178].
The SK Hamiltonian has the same form of Eq. (3), but now the sum runs over all pairs of spins. Also if, for example, the couplings have a Gaussian distribution, one requires that $J=\tilde{J}/N^{\frac{1}{2}}$ in order to make the thermodynamic limit of the model consistent, where $\tilde{J}$ is an intensive number of order one.
This mean-field model has a usual paramagnetic phase for high temperature. The delicate point, which will be the source of trouble and delight, is that, because the coupling are quenched, we have to compute the disorder average not of the partition function $Z$ (which would anneal the disorder degrees of freedom) but of its logarithm, $\ln Z$. That is, we average physical observables in a given disorder realization (a sample) and then we average these thermal expectation values over the disorder. To do this, EA and SK use the identity (the so-called replica trick)
$\ln Z_{J}=\lim_{n\to 0}\frac{Z_{J}^{n}-1}{n}\,$ (7)
and apply it by computing $Z^{n}_{J}$ for integer $n\geq 1$ and eventually taking the limit for $n\longrightarrow 0$. $Z_{J}^{n}$ can be interpreted as the partition function of $n$ copies of the system that are defined in the same disorder landscape (replicas). In principle, this procedure works in the thermodynamic limit if there are no singularities in $n$ when $n\to 0$. In the SK model, however, this approach is successful, in spite of the fact that there is a singularity.
When the temperature decreases below a critical value $T_{\mathrm{g}}$ the correct solution found at high $T$ under the assumption of symmetry between replicas turns out to be undoubtedly wrong. Not only does it not agree with Monte Carlo simulations but, what is far more damning, it develops a negative entropy for low-enough $T$.
Kirkpatrick and Sherrington [178] refined their analysis two years after their first, groundbreaking paper [260]. They described the same situation, but with an analysis based on the assumption of replica symmetry (RS). Before this second SK paper, Thouless et al. [273] had introduced a different approach, avoiding the replica trick, that shed further light on the low-$T$ behavior of the SK model. Their so-called TAP equations, together with their many variations, have been for many years and continue to be important for the analysis of spin-glass physics.
De Almeida and Thouless [93] analyzed the situation and established that the RS solution of the SK model in the high-$T$ (paramagnetic) region was stable, but was unstable at low $T$, with the instability persisting in a magnetic field. Parisi [237] started from here in his (correct) solution of the mean-field spin glass, beginning with the assumption that replica symmetry is broken in the low-temperature phase. Space limits the discussion of his solution. The interested reader is referred to Mézard et al. [213], which includes some of the original work, or to more recent accounts [101, 95, 83].
The above-described mean-field approximation is only correct above the so-called upper critical dimension $D_{\text{u}}$, which for spin glasses is $D_{\text{u}}=6$ [73]. On the other hand, below the lower critical dimension $D_{\text{l}}=2.5$ [124, 197], the low-temperature phase disappears. As a consequence, $D=3$ lies in the non-trivial region. There has been an intense quest for a successful extension of the Parisi theory to $D=3$ in the renormalization-group style [see for example 98, 95]. Because of the infinity of states that appear in Parisi’s solution for the mean-field theory, renormalizing the system is a daunting task. Today, after close to 50 years of studies, there still remains much need for clarification. The interested reader can start from De Dominicis et al. [98] and De Dominicis and Giardina [95], then continue with the following papers and included references: Bray and Moore [68, 69], Bray and Roberts [73], De Dominicis and Kondor [96, 97], Temesvári et al. [269], De Dominicis et al. [94], Temesvári [268].
II.2 Main theoretical pictures for the low-temperature phase
As explained above, the EA model can only be solved in mean field, i.e., in the limit of a large spatial dimension. In $D=3$, therefore, several competing theoretical pictures coexist. This section briefly describes them.
  • Droplet. The droplet picture assumes that the low-temperature phase of spin glasses in the absence of an external magnetic field is composed of only two pure states related by global spin-flip symmetry. The droplet model was developed from a phenomenological theory [113] as well as from the Migdal-Kadanoff approximation of the renormalization group [71]. The behavior of the low-temperature phase derives from the assumed existence of compact excitations (droplets) in the ground state, with a fractal surface. It is also assumed that the energy of these droplets of linear dimension $L$ scales as $L^{y}$. These authors showed that $y\leq(D-1)/2<D-1<D_{\text{s}}<D$, where $D_{\text{s}}$ is the fractal dimension of the surface of the droplets and $D$ is the spatial dimension. Further, the free-energy barriers for the dynamics grow as $L^{\psi}$, with $D-1\geq\psi\geq y$. The presence of an external magnetic field destroys the spin-glass order at low temperatures. The Chaotic Pairs (CP) scenario [223] is similar to the droplet picture with the peculiarity that the two pure states (in the absence of a magnetic field) depend chaotically on the system’s size.
  • Trivial-non-Trivial (TNT). In the droplet model (in the absence of a magnetic field) the probability distribution $P(q)$ of the overlap $q$ between two pure states, and the link overlap, are both trivial (i.e., $P(q)$ is composed of two Dirac delta functions and one Dirac delta function, respectively). In the TNT approach, $P(q)$ exhibits a non-trivial probability distribution, whereas the link overlap exhibits a trivial probability distribution [183, 234].
  • Replica Symmetry Breaking (RSB). In this scenario, which consists in extending the features of the mean-field solution to finite dimensions [203, 95], the spin glass has infinite pure states not related by any kind of symmetry. However, this infinite set of states can be organized in an ultrametric fashion. Furthermore, the excitations of the ground state are space-filling. Finally, in this theoretical approach the spin-glass phase is stable to small magnetic fields (i.e., there exists a phase transition between a paramagnetic phase and a spin-glass phase in the presence of a magnetic field).
Figure 12: The main quantitative difference between the competing theoretical pictures of spin-glass dynamics is related to the value of an exponent $\theta$, the replicon, see Eq. (23) and Marinari et al. [203, 202]. This exponent is non-zero for RSB and zero in the droplet picture, at least in the long-time limit. When represented as a function of a scaling variable $x$ [a properly rescaled $1/\xi(t_{\mathrm{w}},T)$, see 34], the values of $\theta$ obtained for several temperatures fall on a single curve. In addition we show a droplet extrapolation [dashed line, $\theta(x\!\to\!0)\!\propto\!x^{\iota}$, with $\iota\!\approx\!0.15$] and an RSB extrapolation [full line, $\theta(x\!\to\!0)\!>\!0$]. Even the droplet extrapolation, which has $\theta(x)\to 0$ as $x\to 0$, predicts sizable values of $\theta$ for the values of $x$ relevant to experimental length scales (black dots). Data taken from Baity-Jesi et al. [34].
One of the main messages of this review is that the controversy discussed above is of no real consequence for understanding the off-equilibrium dynamics of spin glasses at experimentally relevant scales (see Fig. 12).
II.3 Phenomenological dynamical model on an ultrametric tree
Figure 13: Three representative spin-glass states at a measurement temperature $T_{\mathrm{m}}$ are labeled $\alpha,~{}\beta,~{}{\text{and}}~{}\gamma$. Their overlaps $q$, defined in Eq. (9), satisfy the ultrametric topology $q_{\alpha\gamma}=q_{\beta\gamma}\leq q_{\alpha\beta}$. The self-overlap is independent from the state $q_{\alpha\alpha}=q_{\beta\beta}=q_{\mathrm{EA}}(T)$. When the temperature is lowered from $T_{\mathrm{m}}$ to $T_{\mathrm{m}}-\delta T$, the states “foliate” to daughter states, each of which now possesses the self-overlap $q_{\text{EA}}(T_{\mathrm{m}}-\delta T)$. Figure reproduced from Paga et al. [231].
The Parisi solution of the Sherrington-Kirkpatrick mean-field model is pictured in Fig. 13 in terms of a “tree” that is organized in such a way that ultrametric symmetry relates its states (denoted as $\alpha$, $\beta$,…). Immediately after a temperature quench from above $T_{\mathrm{g}}$ to a measurement temperature $T_{\mathrm{m}}$, the spin-glass states have a “self-overlap” $q_{\alpha\alpha}(T_{\mathrm{m}})\equiv q_{\text{EA}}(T_{\mathrm{m}})$, the Edwards-Anderson order parameter [104],
$q_{\alpha\alpha}\equiv q_{\text{EA}}(T_{\mathrm{m}})={\frac{1}{N}}\sum_{i}\langle\sigma_{i}(t=0)\rangle_{\alpha}^{2},$ (8)
where $\langle\cdot\cdot\cdot\rangle_{\alpha}$ represents a thermal average restricted to a single pure state ($\alpha$).
There are no dynamics associated with pure states in that the free-energy barriers between them are infinite. Experimentally though, time-dependent phenomena are observed. It took a large amount of chutzpah to suggest that between the pure states there could exist states between which the free-energy barriers were finite and even more to suggest that these states between infinite barriers were self-similar: no matter the temperature, the same ultrametricity held for metastable states, which were, then, organized by the same tree as the pure states. The interpretation of dynamical measurements went even further to suggest that the finite barriers between metastable states were temperature dependent, and apparently diverged into the pure states as the temperature was lowered [142], though of course the approach to divergence could not be accessed experimentally.
There is no theoretical proof that these observations are correct. Nevertheless, they have proven remarkably prescient experimentally [283]. A useful picture with finite barriers between metastable states is exhibited in Fig. 14.
Figure 14: Hierarchical organization of metastable states. The coarse-grained free-energy surface is represented at each level corresponding to a given temperature. When temperature is decreased, each valley is subdivided into others. The times $t_{\mathrm{w}}$ and $t_{\mathrm{w}}^{\prime}$ are indicated that are necessary to explore, at $T_{\mathrm{m}}$ and $T_{\mathrm{m}}-\delta T$, respectively, the region of phase space bounded by the same barriers. The closest common ancestor to all states within the space bounded by $\Delta_{1}$ at $T_{\mathrm{m}}$ and $T_{\mathrm{m}}-\delta T$ is the same, and its corresponding value of the overlap function is $q_{1}$. The sketch also shows that, as the system explores more of phase space, it encounters ever-increasing barrier heights, and that the free-energy surface has a self-similar structure. Reproduced from Fig. 6 of Lederman et al. [188].
The dynamics associated with the constructs exhibited in Figs. 13 and 14 follows from their structure. At $t=0$, the temperature is quenched from above $T_{\mathrm{g}}$ in the paramagnetic phase to a measuring temperature $T_{\mathrm{m}}$ in the spin-glass phase. Instantaneously, the spin glass is found in some state $\alpha$ at $T_{\mathrm{m}}$ with self overlap $q_{\alpha\alpha}=q_{\text{\tiny EA}}(T_{\mathrm{m}})$ [Eq. (8)]. As time progresses, the quenched initial state organizes from its paramagnetic structure into a progressively growing spin-glass state, expressed through a growing spin-glass coherence length $\xi_{\mathrm{micro}}(t)$, which is a function of the time elapsed since the quench, termed the waiting time $t_{\mathrm{w}}$. As $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ grows, larger and larger free-energy barrier heights are occupied, having been created by the growing coherence length. The overlap $q_{\alpha\beta}$ with the initial state diminishes and is given by
$q_{\alpha\beta}(t_{\mathrm{w}})={\frac{1}{N}}\sum_{i}\langle\sigma_{i}^{\alpha}(t_{\mathrm{w}}=0)\sigma_{i}^{\beta}(t_{\mathrm{w}})\rangle,$ (9)
where the state has evolved from its initial quenched state $\alpha$ to $\beta$. This is shown in Fig. 13 for the ultrametric tree.
Experimentally, one needs a finite magnetization to make a measurement. This requires a magnetic field that can polarize the otherwise non-magnetic quenched state. Bouchaud [65] introduced another piece of phenomenology, a Zeeman energy $E_{\text{Z}}$, which reduces each of the barrier heights and is given by [159]
$E_{\text{Z}}=-M_{\text{FC}}H=-\chi_{\text{FC}}N_{\text{s}}H^{2},$ (10)
where $\chi_{\text{FC}}$ is the field-cooled magnetization susceptibility per spin when the spin glass is cooled in a magnetic field to $T_{\mathrm{m}}$. $N_{\text{s}}$ is the number of correlated spins within a volume subtended by $\xi(t_{\mathrm{w}})$,
$N_{\text{s}}=\xi(t_{\mathrm{w}})^{D-\theta/2},$ (11)
where $\theta$ is the replicon exponent [33, see also Eq. 23] and $D$ is the spatial dimension. Thus, combining the reduction of the energetic barrier, Eq. (10), to the peak of the relaxation function $S(t,t_{\mathrm{w}};H)$ of Eq. (2) one obtains:
$\ln\left[\frac{t^{\mathrm{eff}}_{\mathrm{H}}}{t^{\mathrm{eff}}_{\mathrm{H\to 0^{+}}}}\right]\propto N_{\text{s}}H^{2}\propto\xi^{D-\theta/2}\,H^{2}\,.$ (12)
It is commonplace to age the spin glass over a time $t_{\mathrm{w}}$, then probe its time dependent magnetization at a time $t$ after a magnetic field is turned on at $t_{\mathrm{w}}$, the so-called zero-field-cooled (ZFC) magnetization. Alternatively, the system can be prepared in the presence of a magnetic field, at a time $t$ after which the field is reduced to zero at time $t_{\mathrm{w}}$, the so-called thermoremanent magnetization (TRM); see Sec. I.2. The spin-glass correlation length is then denoted by $\xi(t,t_{\mathrm{w}};H)$.
Experiments can only extract a spin-glass correlation length through Eq. (12), see Sec. II.6. As a consequence, and because of a plethora of different correlation lengths accessible in simulations, we refer to it as $\xi_{\text{Zeeman}}(t,t_{\mathrm{w}};H)$. It will be shown in Sec. V.2 that it is found to be roughly equivalent to the $\xi_{\text{micro}}(t,t_{\mathrm{w}};H)$ that is extracted through simulations in fixed-temperature protocols. The first use of Eqs. (1012) to extract $\xi_{\text{Zeeman}}(t,t_{\mathrm{w}};H)$ can be found in Joh et al. [159]. A subsequent analysis in terms of the phenomenological model discussed here can be found in Paga et al. [231]. The use of Fig. 14 to display both the self similarity of the metastable states between the infinite barriers of the Parisi pure states, and their evolution toward the pure states can be found in Hammann et al. [142].
II.4 The elusive, but crucial, temperature chaos
Among the many interesting properties of spin glasses, the contextual measurement of both rejuvenation and memory effects has been an astonishing finding [162], which in turn asked for a theoretical understanding, eventually aided by numerical simulations of the EA model. These effects occur in the out-of-equilibrium regime and thus will be mainly covered in Sec. VI.2. Rejuvenation may, however, be related to an equilibrium effect, temperature chaos, which we briefly describe in this section. Furthermore, as we explain in Sec. II.7, temperature chaos is believed to be the main difficulty that impedes reaching thermal equilibrium even for a very small spin-glass sample.
The term chaos refers to the strong sensitivity of the spin-glass state —the one dominating the Gibbs-Boltzmann (GB) measure for $T<T_{\mathrm{g}}$— to the change of some parameter in the GB measure: either the temperature $T$ (temperature chaos), the couplings $\{J_{ij}\}$ (disorder chaos) or the external magnetic field $H$ (field chaos). These effects require the existence of a spin-glass phase, so they cannot take place in any dimensionality. When they are present, they have very different intensities, temperature chaos being by far the weakest. Indeed, for quite a long time the very existence of temperature chaos was controversial [see, e.g., 226, 225, 61, 217, 62, 184, 253, 259, 167, 246, 111], while field chaos and disorder chaos were evident.
In order to quantify the intensity of the chaotic effects, one needs to measure the statistical similarity between two typical configurations generated according to two GB measures differing in a small $\Delta T$ (for temperature chaos) or small $\Delta J$ (for disorder chaos). For example, for temperature chaos, one could measure the statistics of the following chaotic parameter
$X^{J}_{T_{1},T_{2}}=\frac{\langle q^{2}_{T_{1},T_{2}}\rangle_{J}}{(\langle q^{2}_{T_{1},T_{1}}\rangle_{J}\langle q^{2}_{T_{2},T_{2}}\rangle_{J})^{1/2}}\,,$ (13)
where the thermal averages are computed for a given sample indicated by the superscript/subscripts $J$. While the denominator is just a normalization factor to ensure that $X^{J}_{T,T}=1$, the numerator measures the similarity between replicas at temperatures $T_{1}$ and $T_{2}=T_{1}+\Delta T$.
If the sample under consideration is not chaotic, we expect the statistics (e.g., the variance) of overlap $q_{T_{1},T_{2}}$ between replicas at temperatures relatively close to one another to be very similar to those of $q_{T_{1},T_{1}}$ and $q_{T_{2},T_{2}}$, thus giving $X^{J}_{T_{1},T_{2}}\simeq 1$. On the contrary, under chaotic conditions we expect $q_{T_{1},T_{2}}$ to take the smallest possible value (which is zero if the external field is absent). In other words, in a chaotic sample, the configurations at temperatures $T_{1}$ and $T_{2}$ are typically very different (often maximally different).
This is a very relevant physical phenomenon as it would imply that such a chaotic spin-glass sample would develop very different long-range order at temperatures $T_{1}$ and $T_{2}$. If chaos is present, it would be the most natural explanation for rejuvenation: the long-range order developed in the spin glass at the higher temperature $T_{2}$ would be replaced by a new long-range-ordered state when the sample is cooled down to the lower temperature $T_{1}$. How the first ordered state could be eventually recovered when the sample is heated back to $T_{2}$ is discussed in Sec. VI.2 and is the basis of the memory effect.
The approaches to studying and quantifying chaotic effects in spin glasses have followed different paths: a scaling approach based on the droplet model (which is expected to be correct for low-enough dimensionalities); the solution to mean-field models, like the SK model (which is believed to be exact for high-enough dimensionalities); and the approach based on numerical simulations.
According to the droplet theory, the low-temperature properties of a spin glass are controlled by a $T=0$ fixed point, and the large-scale excitations from the ground state (compact droplets with fractal dimension $D_{\text{s}}$) are controlled by an exponent $y$, which is positive (negative) if a spin-glass phase is present (absent). Using scaling arguments, Bray and Moore [71] showed that a perturbation of typical intensity $\Delta J$ to the couplings would destabilize the ground state on length scales larger than $\ell_{\text{c}}\approx(J/\Delta J)^{1/(D_{\text{s}}/2-y)}$, provided $D_{\text{s}}/2-y$ is positive.
Moreover, they extended the argument to a temperature change $\Delta T$ (inside the spin-glass phase), finding again that a typical configuration would be modified on length scales larger than $\ell_{\text{c}}\sim(\Delta T)^{-1/1/(D_{\text{s}}/2-y)}$. So, for any temperature in the spin-glass phase, the overlap $q_{T_{1},T_{2}}$ would eventually trend to zero for large sizes. Within the droplet model, the picture one gets is very simple (probably too simple): the spatial correlation function between two replicas at different temperatures decays exponentially with a characteristic length scale $\ell_{\text{c}}$. Sample-to-sample fluctuations are not discussed, but we shall see that they play an important role in simulations.
A complementary analytical approach can be developed for the mean-field SK model through the replica method [213]. In this case, it is more convenient to compute the free-energy cost $\Delta f(q_{\text{c}})=f(q_{\text{c}},T_{1},T_{2})-f(T_{1})-f(T_{2})$ to keep two replicas at different temperatures $T_{1}$ and $T_{2}$, with an overlap $q_{\text{c}}=q_{T_{1},T_{2}}\neq 0$. Indeed the probability distribution function of that overlap is
$P_{N}(q_{\text{c}})\propto\exp[-N\Delta f(q_{\text{c}})]\;,$ (14)
which becomes a delta function at $q_{\text{c}}=0$ for $N\to\infty$, unless $\Delta f(q_{\text{c}})=0$ in a finite range around $q_{\text{c}}=0$. The free-energy cost $\Delta f(q_{\text{c}})$ is so small that temperature chaos was not found for a long time until the analytical work by Rizzo and Crisanti [253], who carried out a perturbation expansion close to the critical temperature ($T_{1}=T_{\mathrm{g}}-\tau_{1}$ and $T_{2}=T_{\mathrm{g}}-\tau_{2}$) up to the ninth order in $\tau=T_{\mathrm{g}}-T$ to find
$\Delta f(q_{\text{c}})=\frac{12}{35}|q_{c}|^{7}(\tau_{1}-\tau_{2})^{2}\,.$ (15)
Close to the critical point, $q_{\text{c}}\sim(T_{\mathrm{g}}-T)$. See also the numerical results by Billoire and Marinari [62], consistent with these findings.
This analytical result has convincingly clarified that spin glasses do have chaos in temperature (at least in the mean-field approximation), but the intensity of this phenomenon is very weak. This explains why it had not been observed for a long time in numerical simulation of the SK model [277, 263, 220, 64].
The elusiveness of temperature chaos is enhanced when we move away from the realm of mean-field models and study short-range systems. Numerical simulations of the EA model tell a complicated story, where sample-to-sample fluctuations play a key role and which builds on more recent equilibrium studies. Since these 3D results will be directly needed to understand some aspects of the non-equilibrium dynamics, we leave them for Sec. III.1.
II.5 Advent of spin-glass single crystals
Up to recent years, nearly all experiments on spin glasses were done on poly-crystalline samples. A typical example is CuMn dilute alloys, where often the crystallite size is found to be distributed around a characteristic length of approximately $80\,\text{\AA}$ [137, 254]. At first sight, given that poly-crystalline Cu is such a good conductor, one might be inclined to ignore grain boundaries. However, what drives the spin-glass nature in dilute magnetic alloys is the RKKY interaction. Its long-range and oscillatory nature is associated with the sharpness of the Fermi surface.
Figure 15: An as-grown crystal with part of the alumina crucible still attached. A small secondary grain is outlined by a marker. Later, acid etching of the ingot reduced the size and number of the secondary grains, indicating that they are shallow. Reproduced from Fig. 5 of Zhai et al. [299].
However, as Kasuya [166] has shown, electron scattering will change the long-range $\cos(2k_{\text{F}}\,r_{ij})/[k_{\text{F}}\,r_{ij}]^{3}$ dependence of the RKKY interaction to $r_{0}/[(r_{0}^{2}-r_{ij}^{2})^{2}+4r_{0}^{2}r_{ij}^{2}]$, where $k_{\text{F}}$ is the Fermi wave vector, $r_{ij}$ is the distance between spins $i$ and $j$, and $r_{0}$ is a characteristic scattering length that we take to be of the order of the crystallite size. The lack of an oscillatory character removes frustration, a requirement for spin-glass behavior (in addition to randomness), and thus effectively decouples the individual spin-glass grains, limiting the growth of the spin-glass correlation length. Adding to this difficulty is the distribution of grain sizes that must be taken into account for time-dependent properties.
All of the complications associated with poly-crystalline materials can be avoided by the use of single crystals. In that regard, a major step forward took place at the Materials Preparation Center (MPC) of the Ames Laboratory, U.S. Department of Energy, under the direction of D. L. Schlagel. She was able to fabricate two single-crystal CuMn boules with $T_{\mathrm{g}}=31.5$ K and 40.6 K, respectively. The details can be found in Appendix A of Zhai et al. [299]. A picture of the $T_{\mathrm{g}}=$ 31.5 K boule is reproduced in Fig. 15. Laue X-ray diffraction patterns were taken at various points along the length of the body of the same boule. A representative example is exhibited in Fig. 16, confirming that the majority of the body is one single grain. The spots make a clear pattern that is a qualitative measure of crystal quality. There is no evidence of twinning or strain in the crystal.
Figure 16: The Laue x-ray diffraction pattern of the sample in Fig. 15 confirms it is a single crystal; the Cu94Mn6 cube sample was etched in 15% nitric acid. The 6 at. % Mn concentration was estimated from scaling using the observed temperature ($T_{\mathrm{g}}=$ 31.5 K) at which the remanence disappeared. Reproduced from Fig. 6 of Zhai et al. [299].
Use of small samples cut from the 6 at. % CuMn boule allowed growth of the spin-glass correlation length to as much as 150 nm, far larger than any other glassy correlation length measured to date. As such, it enabled observation of a phenomena predicted from simulations of the Janus Collaboration: a dynamic slowing down of the growth rate of the correlation length as the correlation length grows [299]. Many other observations will be reported in this review that were only possible because of the availability of these single crystals.
II.6 Experimental extraction of $\xi(t,t_{\mathrm{w}};H)$
Experiments can measure the time dependence of the magnetization $M(t,t_{\mathrm{w}};H)$ at a fixed temperature. Nordblad et al. [228] introduced the derivative $S(t,t_{\mathrm{w}};H)$ [we rewrite Eq. (2) for the reader’s convenience]
$S(t,t_{\mathrm{w}};H)=\pm{\frac{\text{d}M(t,t_{\mathrm{w}};H)}{\text{d}\ln t}}\,,$ (16)
with the -- for TRM and $+$ for ZFC experiments. This quantity exhibits a maximum at a characteristic time $t_{\mathrm{w}}^{\text{eff}}$ that scales with the waiting time $t_{\mathrm{w}}$. Extraction of $\xi(t,t_{\mathrm{w}};H)$ from measurements of the time at which $S(t,t_{\mathrm{w}};H)$ has a maximum, yielding $t_{\mathrm{w}}^{\text{eff}}$, derives from Vincent et al. [282] using a free-energy “trap model” and Joh et al. [159] using a free-energy barrier model. These barriers or traps are created by the growth of $\xi(t,t_{\mathrm{w}};H)$ and they are invoked to account for the dynamical properties of spin glasses. Further, the deepest trap or the maximum barrier height controls the dynamics because of the exponential growth of occupancy with increasing barrier height. This can be seen from Fig. 14, and is a direct consequence of ultrametricity.
These works noted that the presence of a magnetic field reduces the depth of all the traps or, equivalently, the height of all the energy barriers, by the amount of energy associated with the presence of the magnetic field, or the “Zeeman energy” $E_{\text{Z}}$. They differed in their expressions for $E_{\text{Z}}$. For magnetic fields of laboratory strength the approach of Joh et al. [159] is the more appropriate.
Figure 17: $-M_{\text{TRM}}/H$ against measurement time $t$ (top), with the solid lines from the analytic fit, for CuMn 6 atfor various changes in magnetic field, and for $S(t,t_{\mathrm{w}};H)$ (bottom) calculated by taking the time derivative of the analytic fit (solid line) of $-M_{\text{TRM}}/H$ (top). Reproduced from Fig. 1 of Joh et al. [159].
The relationship between the maximum barrier height $\Delta_{\text{max}}$ and $E_{\text{Z}}$ is
$\Delta_{\text{max}}-N_{\text{s}}\chi_{\text{FC}}H^{2}=k_{\text{B}}T\,\ln t_{\mathrm{w}}^{\text{eff}}-k_{\text{B}}T\,\ln\tau_{0},$ (17)
where $E_{\text{Z}},~{}N_{\text{s}}$, and $\chi_{\text{FC}}$ are defined below Eq. (10), the relationship between $N_{\text{s}}$ and the correlation length $\xi(t_{\mathrm{w}})$ exhibited in Eq. (11), and $\tau_{0}\approx\hbar/k_{\text{B}}T_{\mathrm{g}}$ is an “exchange time.” Thus, by measuring $t_{\mathrm{w}}^{\text{eff}}$ as a function of magnetic field $H$, one can extract $N_{\text{s}}$ and thence $\xi(t,t_{\mathrm{w}};H)$. This is the only vehicle one has available to determine $\xi(t,t_{\mathrm{w}};H)$ from experiment. The usual approaches of neutron diffraction or X-ray scattering measure only two-spin correlations. Unfortunately, for spin glasses, this averages to zero over sample dimensions. Instead, $\xi(t,t_{\mathrm{w}};H)$ derives from a four-spin correlation function [206, see Sec. IV.2].
The first extraction of $\xi(t,t_{\mathrm{w}};H)$ from experiments was performed by Joh et al. [159]. Using Eq. (16), they displayed the time dependence of $S(t,t_{\mathrm{w}};H)$ for different values of $H$ as displayed in Fig. 17. Following Eq. (17), they were able to plot the log $t_{\mathrm{w}}^{\text{eff}}$ vs $H^{2}$, as exhibited in Fig. 18. The break away from linearity with $H^{2}$ has now been ascribed to non-linear effects [231] rather than any breakdown of the relationships in Eq. (17).
Figure 18: A plot of log $t_{\mathrm{w}}^{\text{eff}}$ [equivalently $N_{\text{s}}\chi_{\text{FC}}H^{2}$ from Eq. (6)] vs $H^{2}$ for CuMn 6 at.($T/T_{\mathrm{g}}=0.83,t_{\mathrm{w}}=480$ s) (top) and for the insulating spin-glass CdCr1.7In0.3S4 ($T/T_{\mathrm{g}}=0.21,t_{\mathrm{w}}=3410$ s) at fixed $t_{\mathrm{w}}$ and $T$ (bottom). The dependence is linear for magnetic fields less than 250 and 45 G, respectively. Reproduced from Fig. 2 of Joh et al. [159].
The extraction of $\xi(t,t_{\mathrm{w}};T)$ follows immediately from Eq. (11) once $N_{\text{s}}\chi_{\text{FC}}H^{2}$ has been determined from plots such as Fig. 18. The combination of the use of this quantity, and the availability of single crystals discussed in Sec. II.5, has enabled the first direct comparisons between numerical and experimental measurements of the growth of spin-glass domains.
This fundamental question of relating the experimental correlation length derived from the Zeeman effect, which will be termed $\xi_{\text{Zeeman}}(t,t_{\mathrm{w}};T)$, to the different length scales extracted from simulations (see the glossary in Sec. I.3) will be the topic of Secs. V.2 and V.3.
II.7 Spin-glass simulations and dedicated computers
Simulating the low-temperature behavior of a spin glass is a challenge common to both physicists and computer scientists.
Interest from computer scientists arises because finding the ground state of a spin glass is an archetypically hard computational problem. Indeed, if the interactions of the spin glass can only be drawn in a non-planar graph, [^1] A planar graph is one that can be drawn on a piece of paper without edge crossings, such as the square lattice with open boundary conditions. The difference between planar and two-dimensional is important, as discussed below. then finding the ground state belongs to the category of NP-complete problems [40, 155]. Should a truly efficient algorithm for finding the ground state of a non-planar spin glass be discovered, efficient algorithms for a large category of hard computational problems (such as the traveling salesman problem) would follow [236].
Physicists are often more interested in the behavior of a system at finite temperature $T$ (one could say that physicists wish to minimize the free energy, rather than the energy). However, the physicist and the computer-scientist problems are tightly connected, because the free energy becomes the energy when we approach $T\!=\!0$.
Difficulties faced by physicists depend strongly on whether they consider equilibrium physics or off-equilibrium dynamics:
  • Equilibrium studies are (unfortunately) limited to small systems because reaching thermal equilibrium may require an inordinately long simulation time for some samples [in cubic lattices, equilibration times at low temperatures fluctuate from sample to sample by orders of magnitude, 12, 294, 109]. Further computational difficulties are posed by the fact that a large number of samples need to be simulated in order to get meaningful results, because quantities of interest wildly fluctuate from sample to sample.
  • Non-equilibrium studies are conducted on lattices of large linear size $L$, sufficient to ensure that the coherence length $\xi(t,t_{\mathrm{w}};T)$ is always much smaller than $L$. This condition is required in order to compare with laboratory correlation lengths, which, in general, are smaller than physical sample sizes (with the exception of thin films, see Sec. V.1). Indeed, in space dimension $D$, the system can be approximated as composed of $[L/\xi(t)]^{D}$ more or less independent samples of effective size $L_{\text{eff}}\approx\xi(t)$. Hence, in close analogy to experiments, a simulation of just one very large sample automatically carries out the average over samples. The main difficulty is that one is limited to simulation algorithms that mimic natural dynamics, such as Metropolis [see, e.g., 262] and, as already explained in Sec. II.6, the natural growth of $\xi(t,t_{\mathrm{w}};H)$ is slow. This slowness calls for very long simulation times on few very large lattices.
Simulating spin glasses is not always hard. Efficient algorithms for finding the ground state are known for planar interaction graphs. It is also important to note that planar and two-dimensional are not quite the same thing: the bilayer that was shown in Barahona [40] to pose an NP-hard problem is certainly two-dimensional but non-planar. The square lattice with periodic boundary conditions is mildly non-planar, but still allows one to find ground states for very large lattices [176, 76]. For $T>0$, the problem is also simpler on planar (or quasi-planar) graphs, such as the square lattice with periodic boundary conditions, for which non-local simulation algorithms (named cluster algorithms) can shorten simulation times by orders of magnitude [151, 286].
Unfortunately, the situation is much worse for non-planar graphs, such as the simple-cubic lattice used in most $D=3$ simulations. At $T\!=\!0$, ground states have been found in cubic lattices with size $L\!\leq\!14$ with the help of a heuristic algorithm [199]. See Caracciolo et al. [76] for information on sure solvers. Many algorithms have been proposed for $T\!>\!0$: simulated tempering [198], umbrella sampling [see, e.g., 48, 156], the Wang-Landau algorithm [284] or population annealing [see, e.g., 196]. Nevertheless, the best-performing algorithm for $T\!>\!0$ equilibrium simulations appears to be parallel tempering [129, 152, 270], which, with the help of a custom-built computer, has been able to equilibrate cubic lattices of size $L\!=\!32$ at $T\!\approx\!0.64T_{\mathrm{g}}$ [13]. Temperature chaos, Sec. II.4, is the physical mechanism that makes it difficult for algorithms that exploit the correlation at different temperatures (e.g., simulated annealing or parallel tempering) to reach thermal equilibrium at even lower temperatures [111, 58]. [^2] In a very chaotic sample, thermalization achieved at a higher temperature would be almost irrelevant for sampling at a lower temperature. Thus, the thermalization process would need to restart at the lower temperature.
Lacking better algorithms, the best we can do is to increase computing power. Computers specifically designed for high-energy physics simulations, such as RTN [22] and APE100 [42], have also been quite useful for spin glasses [see Badoni et al. [23], Ciria et al. [86] in the case of RTN and Marinari et al. [206, 205, 201] for APE100]. More recently, graphics processing units (GPUs) have proven to be extremely powerful for simulating quantum spin glasses [50]. Furthermore, a few computers have been specifically designed for the simulation of spin glasses, such as the Ogielski machine [229], SUE [88], and the Janus I [45, 24] and II supercomputers [28].
The great generality of the spin-glass problem uncovered through the computer-science perspective (Sec. I.1) has encouraged people to design and build computers for finding ground states of spin glasses of various types. This includes D-wave’s quantum annealing chips [160, 210], as well as hardware based on a variety of algorithms and/or physical principles [see, e.g., 133, 209, 267, 210, 211].
III The equilibrium picture.
III.1 Some important numerical results in $D=3$
An overview of analytical and numerical results for the equilibrium spin-glass phase is beyond the scope of this review [for a recent account see, e.g., 208, and references therein]. This section discusses a couple of results that will be of use for the interpretation of non-equilibrium data.
As explained in Sec. II.1, numerical studies are conducted on an Ising-like system, the Edwards-Anderson model, whose Hamiltonian is defined in Eq. (3). In $D=3$, this system experiences a continuous phase transition at $T_{\mathrm{g}}=1.1019(29)$. See Baity-Jesi et al. [27] for this value of $T_{\mathrm{g}}$ and for critical exponents.
Figure 19: Probability density function $P(q)$ of the spin overlap at $T=0.703$ for different system sizes. A strictly positive plateau is quickly reached in the $q=0$ sector, even if the position of the peaks is still very noticeably shifting with increasing $L$. The inset shows how $P(0)$ grows roughly linearly with temperature. Data from Alvarez Baños et al. [12].
Perhaps the most straightforward quantitative description of the spin-glass phase comes from the probability density function $P(q)$ for the spin overlap $q$ exhibited in Eq. (5). This is plotted in Fig. 19 for several system sizes and a temperature of $T=0.703\approx 0.64T_{\mathrm{g}}$. Notice the two symmetrical peaks at $q=\pm q_{\text{\tiny EA}}(L)$ and, crucially, that the probability density does not go to zero at $q=0$. Many other studies conducted over the years [47, 200, 205, 169, 46, 288] are in agreement on this point. Should $P(q)$ conserve this shape in the large-$L$ limit, the resulting picture would be consistent with RSB —recall Sec. II.2— but this extrapolation is controversial to this day [214, 297, 59, 60, 57, 287]. The dispute is, however, immaterial for our purposes because, in order to model the behavior of non-equilibrium spin-glass dynamics at finite times, we shall only need the value of $P(q)$ for finite sizes.
A second important feature of the equilibrium spin-glass phase is stochastic stability [186, 239]. In simple terms, if a system with random components is generic, i.e., it behaves as a typical random system and is not drastically changed by a small perturbation, then it is stochastically stable. Several papers have conclusively established that the mean-field spin glass has this property [140, 130, 5]. A numerically testable formulation, relying on overlap equivalence [87, 12], is provided by the following relation
$R_{q^{2}}=\frac{\overline{\langle q^{4}\rangle-\langle q^{2}\rangle^{2}}}{\overline{\langle q^{4}\rangle}-\overline{\langle q^{2}\rangle}^{2}}=\frac{2}{3}\,.$ (18)
Fig. 20 plots this ratio and shows that the value $R_{q^{2}}=\tfrac{2}{3}$ is reached for large $L$ below the critical temperature. Overlap equivalence and stochastic stability are closely related to the ultrametric property of the RSB phase [242, 153, 130, 235], which has been more difficult to observe directly [147, 10, 245]. Stochastic stability allows us to establish a direct quantitative relationship between the off-equilibrium dynamics and the equilibrium properties of a spin-glass system (see Sec. IV.4).
The final key property of the low-temperature phase that we shall need is temperature chaos, introduced in Sec. II.4. The main numerical results have been reported in Fernandez et al. [111] and are summarized here.
Figure 20: The ratio of Eq. (18) in the EA model approaches the expected value $R_{q^{2}}=2/3$ for stochastically stable systems as $L$ grows. Reproduced from Alvarez Baños et al. [12].
Figure 21: Different views on $X_{T_{1},T_{2}}^{J}$, Eq. (13), as a function of $T_{2}$ ($T_{1}\approx 0.64T_{\mathrm{g}}$, the vertical line is $T_{2}=T_{\mathrm{g}}$). Top: for all our system sizes, sample-average of $X_{T_{1},T_{2}}^{J}$. Center: for $L=32$, we show $X_{T_{1},T_{2}}^{J}$ for ten samples $S_{a}$ evenly spaced in a list of growing $\tau_{\text{exp}}$. Bottom: dispersion (i.e., square root of variance over the samples) of $X_{T_{1},T_{2}}^{J}$ for $8\leq L\leq 32$. Reproduced from Fig. 2 of Fernandez et al. [111].
The authors of Fernandez et al. [111] have studied the chaotic parameter $X^{J}_{T_{1},T_{2}}$, defined by Eq. (13), with $T_{1}=0.703\approx 0.64\,T_{\mathrm{g}}$ and $T_{2}$ in a range extending from $T_{1}$ to above $T_{\mathrm{g}}$. A summary of their results is exhibited in Fig. 21. The upper panel reports the sample average of the chaotic parameter, Eq. (13). It decreases with system size, but is far from being zero (the predicted value for a model expected to have complete temperature chaos).
The central panel in Fig. 21 provides deeper insight. The chaotic parameter is shown for ten samples that have different autocorrelation times (a proxy for the difficulty in thermalizing that particular sample). In some samples, the chaotic parameter changes very smoothly, while in others sharp changes are observed over a small temperature range. Such large sample-to-sample fluctuations are quantified in the bottom panel, where they are shown to increase with system size in the spin-glass phase.
The overall picture that comes out from the numerical study in Fernandez et al. [111] is that, for the sizes currently accessible, temperature chaos is a phenomenon only visible in rare samples and controlled by the large-deviation principle
$\text{Probability}[X^{J}_{T_{1},T_{2}}>\varepsilon]=\exp[-L^{D}\Omega_{T_{1},T_{2}}(\varepsilon)]\;.$ (19)
The large-deviation function is well approximated by $\Omega_{T_{1},T_{2}}(\varepsilon)\propto|T_{1}-T_{2}|^{b}\varepsilon^{\beta}$ with $\beta\approx 1.7$. The exponent $b$ can be related to the chaos exponents of Sec. II.4 through a scaling argument: $(D_{\text{s}}/2-y)=D/b$. In this refined scenario, however, the length scale $L_{\text{c}}\sim(\Delta T)^{-b/D}\sim(\Delta T)^{-1/(D_{\text{s}}/2-y)}$ plays a different role with respect to the chaotic length scale $\ell_{\text{c}}$ in the droplet model: the large-deviation principle above implies that chaotic samples will become typical for $L>L_{\text{c}}$.
Nonetheless, the experiments measuring temperature chaos are made in the out-of-equilibrium regime. For this reason, numerical simulations that better reproduce such an effect have been reported in Baity-Jesi et al. [30] and will be discussed in more detail in Sec. VI.1.2.
III.2 Brief overview of results in space dimensions larger than three and the spin-glass transition in a field
We start by describing the behavior of spin glasses in space dimensions larger than three in the absence of an external magnetic field.
The critical behavior of the EA model is well described by an $n$-component $\phi^{3}$ field theory with a tensorial coupling in the $n\to 0$ limit [143]. In six dimensions (the upper critical dimension in zero magnetic field, recall Sec. II.1), the phase transition exhibits mean-field exponents with computable logarithmic corrections [256, 257]. Near and below $D=6$, there should be a phase transition with non-trivial critical exponents [143, 136]. This scenario has been tested in numerical simulations in dimensions four and six with very good agreement [23, 285, 243, 207, 3]. The low-temperature phase in four dimensions follows the RSB predictions. Both the overlap and the link overlap probability distributions are non-trivial.
The behavior of the EA model in the presence of a magnetic field is not as well understood. A first attempt is to try the Wilsonian approach [292, 14]. One begins with the mean-field effective Hamiltonian [73] and introduces fluctuations to determine at which dimension these fluctuations are relevant. This determines the upper critical dimension in a field, $D^{H}_{\text{u}}$.
This upper critical dimension can also be defined as the smallest dimension in which the critical exponents take the mean-field values. However, in contrast with the zero-field case, where the effective Hamiltonian provides $D_{\text{u}}=6$, no stable fixed point at zero coupling has been found in six dimensions for one-loop computations [73, 248]. The two-loop computation finds that the origin (Gaussian fixed point) is unstable, but finds another (non-trivial and maybe nonperturbative) fixed point that could describe the transition in a field  [84, 82].
Furthermore, it is interesting to note that the transition in six dimensions could be quasi-first-order [149]. $D^{H}_{\text{u}}\!=\!8$ has been proposed in a recent computation [17]. Finally, an analysis based on high-temperature series finds signatures of a phase transition for $D\geq 6$ [261].
The droplet model finds that there is no stable spin-glass phase in a magnetic field at any finite dimension [212, 113, 71, 115]. A modified version of the droplet [293] finds a phase transition only for $D>6$ [see also 72, 247].
Numerically, four- and six-dimensional models display a continuous phase transition [11, 3] in a magnetic field. Unfortunately, the estimate of the critical exponents in six dimensions did not allow the determination of whether $D^{H}_{\text{u}}=6$. In addition, numerical simulations in three dimensions have not found a transition in a field [25, 26, 232].
As in zero field, numerical simulations in finite dimensions can be supplemented by numerical simulations in one-dimensional lattices with couplings decaying in an algebraic way (the power law of the decay can be related to the dimensions of regular lattices). The results are contradictory: some studies claim that no spin-glass phase in a field survives below six dimensions [170, 168, 278, 279], while other groups find a phase transition [189, 99].
To summarize the situation, it is widely accepted that the upper critical dimension in the absence of a magnetic field is 6. The behavior below six dimensions agrees with the predictions of the RSB theory, at least for the system sizes that one can equilibrate. Instead, there are still many important open issues regarding the equilibrium behavior in an externally applied magnetic field.
IV Off-equilibrium dynamics: time evolution at no (or very small) magnetic field.
IV.1 Aging in experiments
[…]
This interpretation was accepted until the experiment by Chamberlin et al. [81], who measured the FC magnetization, then cut the field to zero to extract the time dependence of the TRM. They found that the TRM time decay depended on the time spent in the FC state. In other words, though the FC magnetization was itself time independent, its subsequent dynamics after a magnetic field change depended on the time spent in the FC state.
This demonstrated that, though the FC magnetization may have been time independent, it was anything but an equilibrium state. Lundgren et al. [194], using ZFC time-dependent measurements, showed similar dynamics. The two are complimentary according to the superposition principle first enunciated by the Uppsala group [228, 193, 100] [we reproduce Eq. (1) for the reader’s convenience]:
$M_{\text{TRM}}(t,t_{\mathrm{w}})+M_{\text{FC}}(t,t_{\mathrm{w}})=M_{\text{FC}}(0,t_{\mathrm{w}}+t).$ (20)
We now know that this relationship is valid only in the $H\rightarrow 0$ limit [231, see Sec. V.3]. That analysis uses the Hamming distance, Hd, the “distance” between spin-glass states associated with free-energy barriers $\Delta(t,t_{\mathrm{w}};H)$. It is defined in terms of the overlap between two states, $\alpha$ and $\beta$, with $q_{\alpha,\beta}$ as in Eq. (9), where the state has evolved from its initial quenched state $\alpha(t=0,t_{\mathrm{w}}=0;H)$ to $\beta(t,t_{\mathrm{w}};H)$ in a magnetic field $H$:
$\mathrm{Hd}={\frac{1}{2}}(q_{\text{EA}}-q_{\alpha\beta})$ (21)
The free-energy barrier separating states is thus proportional to their distance Hd, as exhibited in Fig. 22. The specific relationship between Hd and barrier heights was first developed by Vertechi and Virasoro [280].
Figure 22: A cartoon illustrating the TRM dynamics associated with the reduction of barrier heights $\Delta(t,t_{\mathrm{w}};H)$ by the Zeeman energy $E_{\text{Z}}$. Top: the system with magnetization $M_{\text{FC}}$, the field-cooled magnetization. Bottom: the system after $H$ was cut to zero with magnetization $M=0$. The arrow represent the decay of $M_{\text{FC}}$ after $H\rightarrow 0$. Reproduced from Fig. 11 of Paga et al. [231]
The decay of $M_{\text{FC}}$ after the magnetic field is cut (e.g., a TRM protocol) is controlled by the states that lie at the highest occupied barrier as displayed in Fig. 22. This is a direct consequence of an exponentially increasing occupancy of states because of ultrametric symmetry, as can be inferred from Fig. 13.
The significance of Fig. 22 is that, in the presence of a magnetic field $H$, the barriers are quenched for values of Hd such that $\Delta(\text{Hd})\leq|E_{\text{Z}}|\equiv\Delta_{E_{\text{Z}}}(\text{Hd})$ with $E_{\text{Z}}$ defined by Eq. (10). Equation (21) can be written alternatively to take into account this cancellation. Take $q_{\alpha\beta}(E_{\text{Z}})=q_{\text{EA}}-q(E_{\text{Z}})$ to be the value of $q_{\alpha\beta}$ at the value of $q$ where $\Delta(q)=|E_{\text{Z}}|$. The probability of finding a value of $q$ larger than $q(E_{\text{Z}})$ is zero. That is, the states between $q_{\text{EA}}$ and $q(E_{\text{Z}})$ have been emptied after the magnetic field $H$ has been cut to zero in a TRM protocol.
[…]
Lederman et al. [188] found explicit evidence for an upward curvature of $\Delta$ as a function of Hd in a AgMn 2.56 at.spin glass. The calculation of Vertechi and Virasoro [280] finds a similar behavior. It is for this reason that Paga et al. [231] concluded that the principle of superposition is only valid in the $H\rightarrow 0$ limit.
IV.2 Correlation functions for non-equilibrium dynamics and the coherence length
The purpose of this section is to introduce the correlation functions that will be instrumental in showing that off-equilibrium dynamics conveys information about equilibrium structures.
Figure 23: Temperature-dependence growth of the coherence length. Left: coherence length $\xi_{12}$($t_{\mathrm{w}}$) vs Monte Carlo time, see Eq. (24), as computed for size $L=256$, for several temperatures [$T_{\mathrm{g}}=1.1019(29)$], data taken from  Fernández and Martín-Mayor [110]. One Monte Carlo step corresponds to 1 ps in physical time. For $T\geq 1.3$, we reach equilibrium. Right: To a first approximation, see Sec. V.2 for more, in the non-equilibrium regime $\xi$ grows with time as a power law $\xi(t_{\mathrm{w}})\sim t_{\mathrm{w}}^{1/z(T)}$. Joined points stand for $T\leq T_{\mathrm{g}}$. Note the constant value $z(T>T_{\mathrm{g}})\approx 6$ (cyan circles).
In general, we shall consider a spin glass that has been quickly cooled from some very high temperature to the working temperature $T_{\mathrm{m}}<T_{\mathrm{g}}$. In computer simulations, the cooling rate can be infinite, so that the starting configuration can be completely disordered. An infinite cooling rate is not experimentally achievable. As explained already several times, the system is then allowed to relax for a time $t_{\mathrm{w}}$. The properties of the system can be studied at the time $t_{\mathrm{w}}$, or at some later time $t_{\mathrm{w}}+t$. Different correlation functions are introduced according to whether we consider only one time, or two.
As the waiting time increases, glassy magnetic domains grow. This growth is monitored using the correlation function [206]
$C_{4}(\boldsymbol{r},t_{\mathrm{w}})=\overline{\frac{1}{N}\sum_{\boldsymbol{x}}\langle q_{\boldsymbol{x}}(t_{\mathrm{w}})q_{\boldsymbol{x}+\boldsymbol{r}}(t_{\mathrm{w}})\rangle}.$ (22)
This quantity is self-averaging, hence for a sample of macroscopic size the disorder average can be omitted. The expected behavior at long distance is
$C_{4}(r\gg 1)=\frac{1}{r^{\theta}}f\bigl{[}r/\xi_{\mathrm{micro}}(t_{\mathrm{w}})\bigr{]}\,,$ (23)
where $\xi_{\mathrm{micro}}$ is the size of the coherent regions (the glassy domains), $\theta$ is the so-called replicon exponent [203, 202] and $f(x)$ is a scaling function that decreases faster than exponentially for large values of its argument. Typically [see 43, 44] $\xi_{\mathrm{micro}}$ is taken as the $k=1$ value of
$\xi_{k,k+1}(t_{\mathrm{w}})={\frac{I_{k+1}(t_{\mathrm{w}})}{I_{k}(t_{\mathrm{w}})}}\,,\ I_{k}(t_{\mathrm{w}})=\int_{0}^{\infty}\mathrm{d}r\,r^{k}C_{4}(r,t_{\mathrm{w}})\,.$ (24)
When applying Eq. (24), care must be taken to check for finite-size effects and to deal with the noisy tails of $C_{4}$. This can be done in a self-consistent manner [see 44, 34]. Fig. 23 exhibits the sluggish growth of $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$.
Figure 24: Dynamic heterogeneities and some non-commuting limits. Top: $C(t,t_{\mathrm{w}})$, see Eq. (25), as a function of $t$ for $t_{\mathrm{w}}=4^{i}$ , $i=3;\dots;16$ (lines, $t_{\mathrm{w}}$ grows from bottom to top). We also plot $\lim_{t_{\mathrm{w}}\to\infty}{C}(t,t_{\mathrm{w}})$ (points) which generates an enveloping curve that ends at $q_{\text{EA}}$ (horizontal line), recall Eq. (26). Bottom: The correlation length $\zeta(C,t_{\mathrm{w}})$, extracted from the correlation function in Eq. (27), as a function of $C^{2}$ (same values of $t_{\mathrm{w}}$ as in the top panel). Note from the top panel that, for fixed $t_{\mathrm{w}}$, $C(t,t_{\mathrm{w}})$ is a monotonically decreasing function of $t$. This length, $\zeta$, indicates the spatial range of the correlations between the changes when the spin configurations at times $t$ and $t_{\mathrm{w}}$ are compared. Figure adapted from Alvarez Baños et al. [13].
Figure 25: The comparison of correlation functions evinces a statics-dynamics equivalence. We plot, with continuous lines, the $C_{2+2}(r,C,t_{\mathrm{w}})$ correlation function of Eq. (27) for several values of $r$, and a fixed $t_{\mathrm{w}}$. $C$ decreases as time $t$ increases, and we show the data for the different values of $t$ as a function of $C^{2}$. This dynamical function coincides with another spatial correlation function $C_{4}(r;L|q)$ (points), computed in equilibrium on a system of size $L$ and conditioned to $q=C$ [see 12, for detailed definitions], if the right equivalence is made between $t_{\mathrm{w}}$ and $L$. In particular, the figure, taken from Alvarez Baños et al. [12], uses a simple statics-dynamics dictionary where non-equilibrium data at time $t_{\mathrm{w}}$ is compared to equilibrium averages for a value of $L$ such that $L=3.7\xi_{12}(t_{\mathrm{w}})$. The main plot shows this comparison for $L=32$, which corresponds to $t_{\mathrm{w}}=2^{31}$, and the inset shows the same for $L=24$, corresponding to $t_{\mathrm{w}}=2^{26}$. The errors in both datasets are comparable and smaller than the point size.
Another correlation function globally checks how much memory is left at time $t_{\mathrm{w}}+t$ of the spin configuration found at time $t_{\mathrm{w}}$:
$C(t,t_{\mathrm{w}})=\overline{\frac{1}{N}\sum_{\boldsymbol{x}}\langle\sigma_{\boldsymbol{x}}(t_{\mathrm{w}})\sigma_{\boldsymbol{x}}(t_{\mathrm{w}}+t)\rangle}\,,$ (25)
It provides, therefore, a direct check of aging [179, 252, 158]. The order of limits is particularly tricky:
$0=\lim_{t_{\mathrm{w}}\to\infty}\lim_{t\to\infty}C(t,t_{\mathrm{w}})\,,\quad q_{\text{\tiny EA}}=\lim_{t\to\infty}\lim_{t_{\mathrm{w}}\to\infty}C(t,t_{\mathrm{w}}),$ (26)
as exemplified in Fig. 24. Sometimes, one changes the variable from $t$ to $C(t,t_{\mathrm{w}})$. This is always possible because $C$ monotonically decreases with $t$ at fixed $t_{\mathrm{w}}$. It is worth stressing that the first limit in Eq. (26) is zero under the assumption of weak ergodicity breaking [89], but a non-zero limit has been recently measured in numerical simulations of spin glasses on Bethe lattices [49].
Finally, one can construct a two-time two-site correlation function that provides spatial information about how the new ordering at later times $t_{\mathrm{w}}+t$ builds by eroding the old ordering pattern at time $t$:
$C_{2+2}(\boldsymbol{r},t,t_{\mathrm{w}})=\overline{\frac{1}{N}\sum_{\boldsymbol{x}}\langle c_{\boldsymbol{x}}(t,t_{\mathrm{w}})c_{\boldsymbol{x}+\boldsymbol{r}}(t,t_{\mathrm{w}})-C^{2}(t,t_{\mathrm{w}})\rangle}\,.$ (27)
where the shorthand $c_{\boldsymbol{x}}(t,t_{\mathrm{w}})=\sigma_{\boldsymbol{x}}(t_{\mathrm{w}})\sigma_{\boldsymbol{x}}(t_{\mathrm{w}}+t)$ is used. Note that $c_{\boldsymbol{x}}(t,t_{\mathrm{w}})=-1$ if the spin at site $\boldsymbol{x}$ has flipped at time $t_{\mathrm{w}}+t$ as compared to time $t$. We name this a defect. The study of these dynamic heterogeneities was initiated in  Castillo et al. [80, 79], Jaubert et al. [157], Aron et al. [20]. Furthermore, Eq. (27) allows the extraction of the correlation length between defects, $\zeta(t,t_{\mathrm{w}})$, which is shown in the lower panel of Fig. 24 by a direct analogy with Eq. (24) [see 44]. By construction, $\zeta(t,t_{\mathrm{w}})$ is upper-bounded by $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$.
An important result that we shall use throughout the rest of this review is that these off-equilibrium quantities can be related quantitatively to their equilibrium counterparts. To do so, one needs to establish a statics-dynamics dictionary, which defines a correspondence between a given finite time $t_{\mathrm{w}}$ (off-equilibrium) and a finite size $L$ (equilibrium). One of the first such dictionaries, based on a single length scale, $\xi_{12}$, is illustrated in Fig. 25. As we shall see below in Sec. IV.4, it is necessary to use several length scales to define a comprehensive statics-dynamics equivalence.
IV.3 Connecting aging with the equilibrium picture: the fluctuation-dissipation relations
To better quantify spin glasses’ aging behavior, we describe the connection between the off-equilibrium aging dynamics and the equilibrium picture with many states and a non-trivial $P(q)$.
The out-of-equilibrium dynamics of strongly disordered systems, like spin glasses, can be solved in a closed form only for models with variables interacting via a fully connected topology. The pioneering work by Cugliandolo and Kurchan [89] considered the fully connected spherical $p$-spin model
$\mathcal{H}=-\sum_{i_{1}<\ldots<i_{p}}^{N}J_{i_{1}\ldots i_{p}}\sigma_{i_{1}}\ldots\sigma_{i_{p}}-\sum_{i=1}^{N}h_{i}\sigma_{i}\;.$ (28)
This model is quite different from realistic spin glasses, as the $\sigma_{i}$ are unbounded real numbers subject to the spherical constraint $\sum_{i=1}^{N}\sigma_{i}^{2}=N$, and interact through all possible groups of $p>2$ spins with very weak couplings in the large-$N$ limit. [^3] The most common choice is to consider Gaussian couplings of zero mean and variance $\overline{J_{i_{1}\ldots i_{p}}^{2}}=p!/(2N^{p-1})$.
The following Langevin equation describes the relaxation dynamics at temperature $T$
$\partial_{t}\sigma_{i}(t)=-\mu(t)\sigma_{i}(t)-\frac{\partial\mathcal{H}}{\partial\sigma_{i}}(t)+\eta_{i}(t)\;,$ (29)
where the first r.h.s. term enforces the spherical constraint and the noise is white $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=2T\delta_{ij}\delta(t-t^{\prime})$.
In addition to the correlation function $C(t,t_{\mathrm{w}})$, Eq. (25), the response function is the key observable:
$R(t,t_{\mathrm{w}})$ $=$ $\frac{1}{N}\sum_{i=1}^{N}\frac{\partial\overline{\langle\sigma_{i}(t+t_{\mathrm{w}})\rangle}}{\partial h_{i}(t_{\mathrm{w}})}\,.$ (30)
Self-consistency equations for $C(t,t_{\mathrm{w}})$ and $R(t,t_{\mathrm{w}})$ can be derived through standard functional methods [89]. Space limitations prohibits writing their equations here, but we can briefly discuss their solution. Following a quench from a high temperature ($T>T_{\text{d}}$) to a low temperature ($T<T_{\text{d}}$), [^4] The dynamical critical temperature $T_{\text{d}}$ is defined by the divergence of the equilibrium correlation time scale. the relaxation dynamics do not reach an equilibrium nor a stationary state on any finite time scale. They remain trapped in an aging regime, where functions of two times depend on both times. In the simplest scenario, we have $C(t,t_{\mathrm{w}})=\mathcal{C}[t_{\mathrm{w}}/(t+t_{\mathrm{w}})]$ and $R(t,t_{\mathrm{w}})=(t+t_{\mathrm{w}})^{-1}\mathcal{R}[t_{\mathrm{w}}/(t+t_{\mathrm{w}})]$, so that relaxation becomes slower while time increases.
Perhaps the most striking result in the Cugliandolo-Kurchan (CK) solution [89] is that, in the aging regime, response and correlation functions are connected by a generalized fluctuation-dissipation relation (GFDR)
$\mathcal{R}(\lambda)=\beta x\;\mathcal{C}^{\prime}(\lambda)\;,$ (31)
where $\lambda=t_{\mathrm{w}}/(t+t_{\mathrm{w}})$ and $x$ is the Parisi parameter in the equilibrium solution to the model. More specifically, close to the equilibrium free energy $f_{0}$, the number of metastable states grows like $\mathcal{N}(f)\sim\exp[N\beta x(f-f_{0})]$, which in turn generates a non-trivial overlap distribution
$P(q)=x\,\delta(q)+(1-x)\,\delta(q-q_{\text{\tiny EA}})\;.$ (32)
Therefore, $x$ is the probability that two replicas at equilibrium are in different states. The connection between the off-equilibrium GFDR and the equilibrium $P(q)$ is foundational to the comprehension of disordered systems from the study of their aging dynamics.
The CK ansatz for the solution to the out-of-equilibrium dynamical equations has been extended by Cugliandolo and Kurchan [90] to the SK model (actually a version of it with soft spins). Even in this more complicated model possessing many diverging time scales [264], the GFDR provides the sought after connection between off-equilibrium dynamics and equilibrium states. Indeed, assuming
$\mathcal{R}(\lambda)=\beta X[\mathcal{C}(\lambda)]\,\mathcal{C}^{\prime}(\lambda)$ (33)
in the aging regime, Cugliandolo and Kurchan [90] solve the dynamical equations and connect the dynamical quantity $X[C]$ to the Parisi distribution of overlaps between states at equilibrium through the relation
$X[\mathcal{C}]=\int_{0}^{\mathcal{C}}\text{d}q\ P(q)\;.$ (34)
Nevertheless, it is fair to remember that the CK solution is based on some crucial hypotheses: weak long-term memory and weak ergodicity breaking. Both are required to decouple the aging dynamics at long times from the initial condition and the fast relaxation at short times. Recently, it has been found that weak ergodicity breaking (which corresponds to the fact that any finite-time configuration is eventually forgotten in the long-time limit) is not always true and must be sometimes replaced by strong ergodicity breaking (meaning that even at very long times the aging dynamics takes place in a restricted space having a positive overlap with the initial condition). Strong ergodicity breaking in off-equilibrium dynamics has been observed in mixed $p$-spin spherical models [117, 118] and in spin-glass models on sparse random graphs [49].
All the above results were obtained for mean-field models. In finite-dimensional models, analytical results are much scarcer. An important one was developed by Franz et al. [122, 123]. These authors considered a generic model with short-range interactions in finite dimensions. The only condition was stochastic stabity, i.e., robustness to small random perturbations (see also Sec. III.1). By studying linear response to the application of long-range perturbations, and assuming that the dynamical averages of one-time quantities (e.g., the energy) converge to their equilibrium values, one can again obtain Eq. (34) relating the equilibrium overlap distribution $P(q)$ to the fluctuation-dissipation ratio
$\lim_{\begin{subarray}{c}t,t_{\mathrm{w}}\to\infty\\ C(t,t_{\mathrm{w}})\to q\end{subarray}}\frac{T\,R(t,t_{\mathrm{w}})}{\partial_{t_{\mathrm{w}}}C(t,t_{\mathrm{w}})}=x(q)\equiv\int_{0}^{q}\text{d}q^{\prime}\ P(q^{\prime})\,.$ (35)
[…]
$T\lim_{\begin{subarray}{c}t,t_{\mathrm{w}}\to\infty\\ C(t,t_{\mathrm{w}})\to q\end{subarray}}\chi(t,t_{\mathrm{w}})=\int_{q}^{1}x(u)\text{d}u\equiv\mathcal{S}(q)\,.$ (36)
Usually, the integrated response $\chi(t,t_{\mathrm{w}})$ is measured as the ratio $m/h$ where $m$ is the magnetization at time $t_{\mathrm{w}}+t$ induced by a field $h$ switched on at time $t_{\mathrm{w}}$. To avoid non-linear effects in the response, methods have been invented that measure $\chi(t,t_{\mathrm{w}})$ without switching on the magnetic field [250, 85].
Figure 26: The integrated response $\chi(t,t_{\mathrm{w}})=m/h_{0}$ (multiplied by the temperature) as a function of the correlation function $C(t,t_{\mathrm{w}})$, measured in the off-equilibrium aging dynamics of the $D=3$ EA model, for different values of $t_{\mathrm{w}}$ and $h_{0}$. It follows closely the dashed curve $S(C)$ obtained via Eqs. (35) and (36) from the equilibrium $P(q)$. This plot, reproduced from Marinari et al. [201], was the first piece of evidence for the validity of statics-dynamics equivalence in spin glasses.
Fig. 26 reproduces the result by Marinari et al. [201], which was the first convincing evidence that the connection between dynamical measurements in the out-of-equilibrium regime and the equilibrium properties, as expressed by Eq. (36), holds in the $D=3$ EA model. Further details on the use of the static-dynamics connection expressed by Eq. (36), and how such a relation needs to be modified in the presence of finite-size effects and growing length scales, will be discussed in Sec. IV.4.
Figure 27: The plot, reproduced from Hérisson and Ocio [148], shows the relation between the integrated response and the correlation measured in experimental spin glasses. It is worth stressing the similarity with the data reported in Fig. 26.
We end this section by presenting the first and most important measurement of the fluctuation-dissipation ratio in experimental spin glasses. Hérisson and Ocio [148] obtained an impressive result by measuring, on the same insulating spin glass at a temperature $T=0.8\,T_{\mathrm{g}}$, both the integrated response $\chi(t,t_{\mathrm{w}})$ and the magnetization autocorrelation $C(t,t_{\mathrm{w}})$. By plotting both quantities parametrically in $t$ for several values of $t_{\mathrm{w}}$ (see Fig. 27), Hérisson and Ocio [148] reported a non-trivial fluctuation-dissipation ratio in experimental spin glasses very similar to the one that was already exhibited in numerical simulations of spin-glass models [202, see, e.g., Fig. 26].
The reader will note, however, that the experimental results in Fig. 27 display a clear time dependence. In other words, the experiment of Hérisson and Ocio [148] was unable to reach the infinite-time limit of Eq. (36). [^5] Herisson and Ocio were well aware of this fact and attempted to extrapolate their results to infinite time. Because very large times mean very large coherence lengths, the sluggish growth of $\xi_{\mathrm{micro}}$ at low temperatures depicted in Fig. 23 suggests that it is highly unlikely that any near-term experiments will be able to reach the limit required in Eq. (36). This raises the question: how should Eq. (36) be used when the limit of large times is unreachable? This is, precisely, the question that we shall address in the next section.
IV.4 The statics-dynamics dictionary
We discuss here the important issue of the quantitative connection between the non-equilibrium evolution of a spin glass (dynamics) and its equilibrium behavior (statics), following Baity-Jesi et al. [36]. Accurate numerical computations of the non-equilibrium fluctuation-dissipation ratio for the EA model have been the basis for establishing a statics–dynamics dictionary. This is of great help for computing controlled extrapolations.
A framework has been emerging in the past few years for theoretical explanations of what can be, and is, observed under actual experimental conditions. One needs to reach a theoretical understanding of what can be measured under the “best” experimental conditions. For example, it was found in the seminal experiments of Joh et al. [159] that the largest spin-glass correlation length one could reach in the laboratory was of order $50~{}a_{0}$, with $a_{0}$ the mean distance between magnetic moments. This was in 1998. Today, because of the availability of single crystals, one can do better, reaching $234~{}a_{0}$ [299, recall Sec. II.5]. The two settings, numerical simulations and experiments, are plagued by different limitations. If they can be overcome, a joint understanding becomes possible.
The root for the theoretical analysis lies in the work of Cugliandolo and Kurchan [89], see also Franz and Mézard [120, 121], which has been discussed in detail in Sec. IV.3. Interesting work attempting to extend their analysis to finite times can also be found in Barrat and Berthier [41], Berthier et al. [55], Belletti et al. [43], Alvarez Baños et al. [12, 13].
The basis for our discussion is the very slow growth of the coherence length in the low-temperature phase of spin glasses (already discussed in detail in Sec. IV.2). Another crucial ingredient for this analysis is the stochastic stability discussed in Sec. III.1 and Sec. IV.3 because it allows the establishment of a quantitative relationship between dynamics and statics, a major issue. As discussed in Sec. IV.3, one can use stochastic stability to derive the GFDR [122, 201].
We shall consider, instead, quantities that can be computed on a lattice of size $L$, like the equilibrium (static)
$x(C,L)\equiv\int_{0}^{C}\text{d}q\;P_{L}\left(q\right)\equiv-\frac{\text{d}\mathcal{S}\left(C,L\right)}{\text{d}C}\,$ (37)
(recall Sec. III.1 and the finite size $P_{L}(q)$ shown in Fig. 19), the dynamic linear response function measured in off-equilibrium experiments by switching on a field $\delta H$ at time $t_{\mathrm{w}}$
$\chi_{L}\left(t,t_{\mathrm{w}}\right)\equiv\lim_{\delta H\rightarrow 0}\;\frac{\delta m(t+t_{\mathrm{w}})}{\delta H}\;,$ (38)
and the corresponding correlation function $C_{L}(t,t_{\mathrm{w}})$ defined in Eq. (25).
We now go to the core issue: how an effective and well-functioning statics-dynamics dictionary (SDD), based on a numerical GFDR, can be generated [36]. It turns out to be crucial that the SDD be based on matching observables, both at the waiting time $t_{\mathrm{w}}$ (where the experimental conditions are set) and at the probing time $t_{\mathrm{w}}+t$ (where experimental measurements are taken).
Numerical simulations of a $3D$ EA spin glass on a system of linear size $L$ and volume $V=L^{3}$ can produce data that follows the relevant experimental protocol. One cools the simulated system from high temperatures $T$ down to, say, $T=0.64\,T_{g}$ in zero magnetic field, $H=0$. The system then evolves during the waiting time $t_{\mathrm{w}}$, generating a growing coherence length $\xi(t,t_{\mathrm{w}};H\!=\!0)$. At $t_{\mathrm{w}}$, one switches on a magnetic field $H$ and evaluates for further times the linear response function, $\chi_{L}(t,t_{\mathrm{w}})$, together with the correlation function $C_{L}(t,t_{\mathrm{w}})$. We denote the infinite-volume limit of different observables by omitting the subscript $L$. For $T<T_{\mathrm{g}}$ the GFDR in Eq. (36) has the form
$\lim_{t,t_{\mathrm{w}}\to\infty}\lim_{L\to\infty}T\chi_{L}\left(t,t_{\mathrm{w}}\right)=\lim_{t,t_{\mathrm{w}}\to\infty}\lim_{L\to\infty}\mathcal{S}\left(C_{L}(t,t_{\mathrm{w}}),L\right)\,,$ (39)
where one must be careful with the order of the limits. The measurement time $t$ has to be scaled appropriately for increasing $t_{\mathrm{w}}$ as in Eq. (36).
Figure 28: The “naive” scaling does not work when the correlation function becomes too small. It also does not work for short times when close to $q_{\text{\tiny EA}}$. Reproduced from Fig. 2 of Baity-Jesi et al. [36].
Taking the limit of large $L$ in the l.h.s. of Eq. (39) means keeping $L\gg\xi_{\mathrm{micro}}$ for the whole time of the experiment (or the simulation). However, in a numerical simulation, the order of the limits in the r.h.s. of Eq. (39), must be reversed. We reach equilibrium (i.e., infinite time) by using some unphysical dynamics (recall Sec. II.7) in a relatively small system. The function $\mathcal{S}\left[q=C_{L}(t,t_{\mathrm{w}}),L\right]$ is computed numerically by integrating the finite size probability distribution for the overlap $P_{L}(q)$. In other words, one uses numerical simulations to mimic experiments by relating the non-equilibrium response of a very large system at finite times to the equilibrium overlap for a much smaller system. In particular, one may find the effective equilibrium size $L_{\text{eff}}(t,t_{\mathrm{w}})$ as the value of $L$ such that $T\chi(t,t_{\mathrm{w}})=\mathcal{S}\left(q=C(t,t_{\mathrm{w}}),L\right)$. [^6] Recall that, if we keep $t$ and $t_{\mathrm{w}}$ fixed, the limit of infinite size can be reached for both $C$ and $\chi$ without much pain in the off-equilibrium experiment (or simulation). Unfortunately, the limit of large $L$ for $\mathcal{S}(q,L)$ is a lot more difficult to obtain.
Figure 29: The improved scaling, which depends on two coherence lengths, see Eq. (41), works well in all measured regimes. Reproduced from Fig. 4 of Baity-Jesi et al. [36].
In a first approach, one tries a single-time ansatz for the effective equilibrium size $L_{\text{eff}}(t,t_{\mathrm{w}})$. For instance,
$L_{\text{eff}}(t,t_{\mathrm{w}})=k\,\xi(t_{\mathrm{w}})\,,$ (40)
This is the method illustrated in Fig. 25, where $k=3.7$ was shown to work to scale the $C_{2+2}$ correlation function. When studying the response function, however, this simple statics-dynamics dictionary does not work, as shown in Fig. 28, where it is represented by the dotted lines. Even allowing a different value of $k$ for each value of $t_{\mathrm{w}}$ (solid lines) does not work. Again, both for short and for long times, we do not get a good description of the data. Our scaling hypothesis was, somehow, too naive.
It is not difficult to understand the discrepancy at large $t$. We expect $\xi$ to grow without bounds in the spin-glass phase, resulting in $\xi(t+t_{\mathrm{w}})$ becoming eventually far larger and very different from $\xi(t_{\mathrm{w}})$.
It is clear at this point that the effective length ${L}_{\text{eff}}$ must be analyzed by considering as relevant both coherence lengths $\xi(t_{\mathrm{w}})$ and $\xi(t+t_{\mathrm{w}})$. The numerical data suggests that, to good accuracy, $L_{\text{eff}}/\xi(t+t_{\mathrm{w}})$ is a function of $\xi(t+t_{\mathrm{w}})/\xi(t_{\mathrm{w}})$. So a new scaling law, depending on two coherence lengths, can be analyzed by assuming
$L_{\text{eff}}(t,t_{\mathrm{w}})=\xi(t+t_{\mathrm{w}})\,h\left[\xi(t+t_{\mathrm{w}})/\xi(t_{\mathrm{w}})\right]\,,$ (41)
with a scaling function of the form $h(x)=k_{1}+k_{2}x^{-c}$. Using numerical data to determine plausible best values for the fitting parameters, one finds that it is legitimate and very sensible to rewrite $h(x)$ in the simpler form
$h\left[\xi(t+t_{\mathrm{w}})/\xi(t_{\mathrm{w}})\right]=k_{1}\left(1+\sqrt{\frac{t_{\mathrm{w}}}{t+t_{\mathrm{w}}}}\right)\;,$ (42)
which depends on a single free parameter. The success of this assumption is clear from Fig. 29, where one sees that the scaling (41) works nicely over all time regimes.
The line of reasoning presented here is based on the idea that the experimental aging response functions contain information about the Parisi order parameter $P(q)$, and is motivated by the fact that the extrapolation to infinite waiting times is cumbersome and sometimes mysterious. The scaling hypotheses (41) and (42), based on assuming that two different length and time scales have to be incorporated in the SDD, allow an orderly, systematic and fully successful extrapolation. When $\xi(t+t_{\mathrm{w}})\gg\xi(t_{\mathrm{w}})$ this dictionary disagrees with the simple SDD base on a single length scale, and correctly reports that, under these conditions, $L_{\text{eff}}\sim\xi(t+t_{\mathrm{w}})$.
V Off-equilibrium dynamics: non-linear response at a single temperature
V.1 Experiments on films and the lower critical dimension
The use of thin-film spin glasses has generated some remarkable dynamical properties. At first glance, it would seem improbable that a sufficient number of spins were contained in a thin film of thickness, say, 2 or 3 nm, for measurement with even the most sensitive of instruments. The situation changed with the introduction of thin-film multilayers comprised of a thin spin-glass layer separated by a non-magnetic layer sufficiently thick to eliminate any communication between the spin-glass layers. These were pioneered by Kenning et al. [175, 172], who used thin layers of CuMn separated by thick Cu layers. The use of the same metal for solvent and separator eliminated the problem of strains at the interfaces.
An example of their data is exhibited in Fig. 30. The peak in the ZFC susceptibility, and its separation from the FC magnetization, moves to lower temperatures as the thickness of the CuMn films is reduced. This “freezing temperature”, $T_{\text{f}}$, is distinguished from the bulk spin-glass condensation temperature $T_{\mathrm{g}}$. It arises from the limitation of the growth of the correlation length by the thickness of the film, thereby never reaching the value required for a true phase transition. Instead, the maximum barrier height, $\Delta_{\text{max}}({\mathcal{L}})$, set by the thin-film parameters, causes an apparent freezing when the laboratory time scale, $\tau_{\text{expt}}$, satisfies
${\frac{1}{\tau_{\text{expt}}}}\approx{\frac{1}{\tau_{0}}}{\text{exp}}[-\Delta_{\text{max}}({\mathcal{L}})/k_{\text{B}}T_{\text{f}}].$ (43)
The dynamics of the CuMn multilayers were first considered by Sandlund et al. [258]. They measured the temperature dependence of the dynamic susceptibility at different observation times from ZFC data for a 3 nm film, exhibited in Fig. 31. Their data were analyzed by Guchhait et al. [137] in terms of a temperature-dependent maximum barrier height [142]. This assumed a hierarchical structure of free-energy states as displayed in Fig. 14. Work by Guchhait et al. [137] also assumed that the coherent volume reached a steady-state maximum when $\xi(t,T)=\mathcal{L}$, the thickness of the CuMn layer.
In order to account for the thin film geometry, Guchhait et al. [137] described the coherent volume as “pancake-like,” with a perpendicular length scale $\xi^{\perp}(t,T)$ and a parallel length scale $\xi^{\parallel}(t,T)$. They defined an effective correlation length,
$\xi^{\text{eff}}(t,T)=\{\xi^{\perp}(t,T)[\xi^{\parallel}(t,T)]^{2}\}^{1/3}.$ (44)
They assumed that, upon nucleation, the correlation length grew as in a bulk sample until it reached the film thickness, then saturated at $\mathcal{L}$, defining a crossover time,
$\xi^{\perp}(t_{\text{co}},T)=\mathcal{L}.$ (45)
The parallel component was more troublesome. After crossover, it was assumed that $\xi^{\parallel}(t,T)$ reflected the behavior of a $D=2$ spin glass for which $T_{\mathrm{g}}=0$, the lower critical dimension for an Ising spin glass being $D_{\text{l}}=2.5$ (Sec. II.1). This was supported by an experimental observation on a thin film of an amorphous GeMn alloy [138] that found 2<$D_{\text{l}}$<3. The authors further assumed that, after $t_{\text{co}}$, $\xi^{\parallel}(t\geq t_{\text{co}},T)$ rapidly equilibrated. Thus, they took
$\xi^{\parallel}(t>t_{\text{co}},T)=\xi_{\text{eq}}^{\parallel}={\mathcal{L}}\bigg{(}{\frac{T_{\text{q}}}{T}}\bigg{)}^{\nu},$ (46)
where $T_{\text{q}}$ is the quenched temperature of Fig. 31 and $\nu$ is the Ising critical exponent of a pure $3D$ model.
Figure 30: Magnetic susceptibility as a function of temperature for Cu0.86Mn0.14/Cu multilayer systems. The open symbols are for ZFC data and the solid symbols are for FC data. The arrows indicated the transition temperatures. The thicknesses are reported in Angstroms. Reproduced from Fig. 14 of Kenning et al. [172].
Figure 31: Temperature dependence of the dynamic susceptibility for a CuMn 13.5 at.multilayer at different observation times. Reproduced from Fig. 1 of Sandlund et al. [258].
Putting all this together, Guchhait et al. [137] hypothesized an effective correlation length in equilibrium after crossover to be given by,
$\xi^{\text{eff}}(t>t_{\text{co}},T)=\mathcal{L}\bigg{(}{\frac{T_{\text{q}}}{T}}\bigg{)}^{2\nu/3}.$ (47)
Simulations (see Sec. V.2), however, have found that the parallel correlation length continues to grow after $t_{\text{co}}$. To probe this time regime would require measurements over time scales much larger than heretofore attained in the laboratory.
V.2 Relating the numerical coherence length $\xi_{\mathrm{micro}}$ and the experimental $\xi_{\mathrm{Zeeman}}$
Figure 32: The growth of the coherence length $\xi_{\mathrm{micro}}$, computed using the microscopic correlation function $C_{4}(r,t_{\mathrm{w}})$, see Sec. IV.2, is compared to the macroscopic $\xi_{\text{macro}}$, which is calculated by reproducing in simulations the experimental recipe to measure $\xi_{\mathrm{Zeeman}}$. The figure shows two alternatives, either the version of Eq. (12), based on a quadratic dependence of the Zeeman energy on $H$, or that of Eq. (48), based on a linear dependence. The microscopic time scale $\tau_{0}=1$ corresponds to a single lattice sweep. The scaling variable $T\ln(t_{\mathrm{w}}/\tau_{0})$ is standard in the experimental literature [219]. Figure taken from Baity-Jesi et al. [33].
Throughout this review, recall the glossary in Sec. I.3, we have considered several length scales in parallel. On the one hand, from theory and simulations, we have the coherence length $\xi_{\mathrm{micro}}$ extracted from the $C_{4}$ correlation function as explained in Sec. IV.2. Such microscopic detail is not accessible to experiments, where instead a $\xi_{\mathrm{Zeeman}}$ based on the Zeeman energy is measured, as described in Secs. 13 and II.6. On the other hand, while simulations have reached longer times with the advent of special-purpose computers, they are still not quite at the experimental scale. This section will show how to relate experimental and numerical length scales and, hence, move towards the joint quantitative analysis of experiment and simulation that will be treated in the following sections.
The first major step in this direction was taken by Baity-Jesi et al. [36] who measured the response function $S$ of Eq. (2), the basic tool of experiments, in simulations of the EA model. This procedure allowed the calculation of a numerical $\xi_{\text{macro}}$ that reproduced the experiments from which $\xi_{\mathrm{Zeeman}}$ was extracted.
The key formula is Eq. (12), where a quadratic dependence of the Zeeman energy on the magnetic field was assumed. One could also compute $\xi_{\text{macro}}$ assuming, instead, a linear dependence [52]. Under that assumption, Eq. (12) would be modified to
$\ln\left[\frac{t^{\mathrm{eff}}_{\mathrm{H}}}{t^{\mathrm{eff}}_{\mathrm{H\to 0^{+}}}}\right]=A^{\prime}\sqrt{N_{f}(t_{\mathrm{w}})H^{2}}\,.$ (48)
and then $\xi_{\text{macro}}(t_{\mathrm{w}})=[N_{f}(t_{\mathrm{w}})]^{1/(D-\theta/2)}$.
The comparison between $\xi_{\text{micro}}$ and the two determinations of $\xi_{\text{macro}}$ using quadratic and linear fits to the magnetic field is exhibited in Fig. 32. The behavior of these three lengths is quite similar.
The results of Baity-Jesi et al. [36] suggested that the numerical $\xi_{\mathrm{micro}}$ would have the same physics as the experimental $\xi_{\mathrm{Zeeman}}$. A major quantitative discrepancy, however, still remained: the aging rates measured in simulations and experiments greatly differed.
The growth of the coherence length $\xi_{\mathrm{micro}}(t_{\mathrm{w}},T)$ was usually expressed through a power law that defined a temperature-dependent dynamic critical exponent $z(T)$,
$\xi_{\mathrm{micro}}(t_{\mathrm{w}},T)\simeq A(T)\,t_{\mathrm{w}}^{1/z(T)}\,,$ (49)
with $z(T)\simeq z_{\text{c}}T_{\mathrm{g}}/T$ and $z_{\text{c}}\equiv z(T_{\mathrm{g}})$. According to the previous discussion, $z(T)$ should be the same for the experimental $\xi_{\mathrm{Zeeman}}$.
Figure 33: Extrapolation of the $\xi_{\mathrm{micro}}$-dependent aging rate $z(T,\xi_{\mathrm{micro}})$ of Eqs. (50) and (51) to the experimental range of values of the coherence length in thin CuMn films. The resulting ratio $Z_{\text{c}}(T)=z(T,\xi_{\mathrm{micro}})T/T_{\mathrm{g}}$ exhibits a very good agreement to the experimental value $Z_{\text{c}}\approx 9.62$ [298], indicated by the straight line running along the experimental temperature range. Critical effects can only be detected for $T>0.7$. Data from Baity-Jesi et al. [34].
Unfortunately, experiments on CuMn films [298] reported an almost constant value of the ratio $z(T)T/T_{\mathrm{g}}\approx 9.62$ [298] whereas numerical simulations found $z_{\text{c}}=6.86(16)$ [43] and $6.80(15)$ [192].
Taking advantage of much more precise measurements of $C_{4}$, the authors of Baity-Jesi et al. [34] were able to find the missing piece of the puzzle by observing that the dynamical critical exponent is actually dependent on $\xi_{\mathrm{micro}}$. Eq. (49) is only an approximation that works for a limited time range and, to make a connection to experimental time scales, one has to consider
$\ln t_{\mathrm{w}}=D(T)+z_{\infty}(T)\,\ln\xi_{\mathrm{micro}}+E(T)\,\xi_{\mathrm{micro}}^{-\omega}\,,$ (50)
where the finite-$\xi_{\mathrm{micro}}$ dependence of $z$ is controlled by an exponent $\omega$. At the critical point, $\omega$ should be the exponent of the first correction to scaling [see 27]. In the spin-glass phase, it would correspond to the replicon exponent, $\omega=\theta$.
Using Eq. (50) we can define an effective dynamic critical exponent (or aging rate) for each value of $\xi_{\mathrm{micro}}$:
$z(T,\xi_{\mathrm{micro}})=\frac{\text{d}\ln t_{\text{w}}}{\text{d}\ln\xi_{\mathrm{micro}}}\,,$ (51)
Eq. (51) can be extrapolated for typical experimental values in thin films to find $Z_{\text{c}}(T)=z(T,\xi_{\text{film}})T/T_{\mathrm{g}}$. These extrapolations are shown in Fig. 33. The agreement with the experimental value is very good. A modification of Eq. (49) based on activated dynamics [66, 53] has less success in reproducing the nearly constant $Z_{\text{c}}(T)$ seen in experiments [34].
We are now in a position to compare simulations with experiments, beginning with the analyses of thin-film dynamics of Sec. IV.1. This step was taken in Fernandez et al. [106], finding a scenario with some similarities, but also some important differences with the experimental observations.
Fig. 34 exhibits the general behavior of $\xi_{\mathrm{micro}}^{\parallel}$ [computed, as usual, with the $k=1$ value of Eq. (24) as applied to the parallel correlation function] as a function of time, for three different temperatures (the critical temperature of bulk systems, and two temperatures below) and for four values of the perpendicular size of the sample $L_{z}$, much smaller than the parallel size.
Figure 34: The longitudinal coherence length $\xi_{\mathrm{micro}}^{\parallel}(T,t)$, as computed in films of thickness $L_{z}$, against the waiting time $t$ after a quench to temperature $T$, for $T=0.98$ (main), $T=1.1$ (upper insert), and $T=0.7$ (lower insert). The critical temperature is $T_{\mathrm{g}}=1.102(3)$ [27]. Also shown, as a reference, purely 3D dynamics [29] and fits to 2D dynamics $\xi_{\mathrm{micro}}^{\parallel}(L_{z},T,t)\approx b(L_{z},T)+a(L_{z},T)t^{1/z_{2D}}$, with $z_{2D}=7.14$ [108], being $b(L_{z},T)$ and $a(L_{z},T)$ the fit parameters.
Instead of the two time regimes invoked by experiment [137], simulations find four.
In the first regime, for small times, the growth of $\xi_{\mathrm{micro}}^{\parallel}$ is indistinguishable from what happens in 3D, consistent with the experimental interpretation for $\xi^{\perp}(t,T)$. However, eventually the growth rate changes (for example for $T=0.98$ and $L_{z}=16$ at a time larger than $10^{4}$) and the system enters a second regime where $\xi_{\mathrm{micro}}^{\parallel}$ grows faster than in $D=3$. After a transient period, in a third regime, $\xi_{\mathrm{micro}}^{\parallel}$ grows as in 2D which, for $T<T_{\mathrm{g}}$, is faster than 3D growth. Finally, the fourth regime corresponds to the saturation of $\xi_{\mathrm{micro}}^{\parallel}$ to its equilibrium value. The fourth regime is completed in the numerical data for $L_{z}=4$ at $T=0.98$, and for all simulated $L_{z}$ at $T=1.1$.
A major difference between the experimental interpretation and the findings of simulations is the rate of growth of the correlation length after crossover. The interpretation of experiments is that $\xi^{\parallel}(t,T)$ is in $D=2$ equilibrium after crossover. This is in opposition to the findings of simulations that $\xi_{\mathrm{micro}}^{\parallel}(t_{\mathrm{w}},T)$ increases even more rapidly than in 3D after crossover. There is clearly need for new experiments to probe thin-film dynamics in this unexplored regime (see Sec. VII.2).
V.3 Direct comparison of CuMn experiments with Janus II simulations
The advent of dedicated supercomputers, such as Janus II, and of single crystals of CuMn [299] opened up the possibility of combining numerical and experimental results for the first time. Given this opportunity, the initial challenge was to define observables accessible to both theory and laboratory. On the one hand, one has direct access to microscopic configurations through simulations. On the other hand, experiments measure the system’s response to external magnetic field perturbations, either switching it on (ZFC) or off (TRM). The previous section showed how these two approaches could be reconciled, by relating the experimental correlation length $\xi_{\mathrm{Zeeman}}$ to the microscopic coherence length $\xi_{\mathrm{micro}}$ measured in simulations. The key was the attempt by Baity-Jesi et al. [33] to measure experimental response functions in simulations. That approach, however, relied on an approximate scaling law, valid only for some range of values of the external field and which broke down close to $T_{\mathrm{g}}$ (Fig. 33), which is precisely the most interesting experimental regime for glassy systems (because it allows the observation of larger coherence lengths).
A more sophisticated treatment was clearly needed. In 2020, a collaboration between experimental and theoretical groups [302, 232] took the next step, beginning with a more precise replication of the experimental ZFC protocol in simulation. Two tricks were required:
  • Because $H=1$ for the EA model with binary couplings corresponds to $\sim 5\times 10^{4}\,\mathrm{Oe}$ [33, 21], to match the experimental scales a dimensional analysis [116] was used to relate $H$ and the reduced temperature $\hat{t}=(T_{\mathrm{g}}-T)/T_{\mathrm{g}}$:
    $\hat{t}_{\mathrm{num}}\approx\hat{t}_{\mathrm{exp}}\bigg{(}\frac{H_{\mathrm{num}}}{H_{\mathrm{exp}}}\bigg{)}^{\frac{4}{\nu(5-\eta)}},$ (52)
    where $\nu=2.56(4)$ and $\eta=-0.390(4)$ are the critical exponents at $H=0$ [27].
  • Experiments define an effective response time $t^{\text{eff}}_{H}$ from the the time at which the relaxation function, $S(t,t_{\mathrm{w}};H)$, peaks. A different strategy was proposed for simulations, where the time at which $S(t,t_{\mathrm{w}};H)$ peaks was found to occur at a constant value of the correlation function $C(t,t_{\mathrm{w}};H)$ of Eq. (25). This allowed a huge simplification for the extraction of $t^{\mathrm{eff}}_{H}$ because
    $C(t^{\mathrm{eff}}_{H},t_{\mathrm{w}};H)=C_{\mathrm{peak}}\,,$ (53)
    and could be solved at $H=0$ as well. Moreover, the scaling law that appears if one displays $\ln(t/t^{\mathrm{eff}}_{H})$ as a function of the Hamming distance $\mathrm{Hd}$, defined in Eq. (21), guarantees that the extraction of $t^{\mathrm{eff}}_{H}$ is not sensitive to the precise value of $C_{\mathrm{peak}}$ [231]:
    $\ln\{t/t^{\mathrm{eff}}_{H}[C_{\mathrm{peak}}(t_{\mathrm{w}})]\}=\mathcal{F}[C(t,t_{\mathrm{w}};H),t_{\mathrm{w}}]\;.$ (54)
    See Fig. 35 for details.
Figure 35: Behavior of the relaxation function, $S(t,t_{\mathrm{w}};H)$ as a function of the correlation function, $C(t,t_{\mathrm{w}};H)$ (left) and as a function of time $t$ (right) for several magnetic fields. The peak’s position occurs at a constant value $C_{\mathrm{peak}}$. The insert shows the scaling law that emerges if time is rescaled by the extracted effective time $t^{\text{eff}}_{H}$ as a function of the Hamming distance $\mathrm{Hd}$. The dashed line refers to the value of $C_{\mathrm{peak}}$. See Zhai et al. [302], Paga et al. [232] for more details.
Because of these observations and tricks, Zhai et al. [302] and Paga et al. [232] were able to introduce a scaling law that describes the system’s response over the entire natural range of variation, allowing a bridge to be built between simulations and experiments:
$\ln\frac{t^{\mathrm{eff}}_{H}}{t^{\mathrm{eff}}_{H\to 0^{+}}}=\frac{\hat{S}}{T}\xi^{D-\theta/2}H^{2}+\xi^{-\theta/2}\mathcal{G}(\xi^{D-\theta/2}H^{2};T)\,.$ (55)
Here $\xi$ stands for $\xi_{\mathrm{micro}}(t_{\mathrm{w}},T)$ [which it is equivalent to $\xi_{\mathrm{Zeeman}}(t_{\mathrm{w}},T)$, recall Fig. 32], $\hat{S}$ is a constant, $D=3$ is the spatial dimension and $\theta$ stands for the replicon exponent $\theta(\tilde{x})$, where $\tilde{x}=\ell_{\text{J}}(T)/\xi(t_{\mathrm{w}},T)$ and $\ell_{\text{J}}(T)$ is the Josephson length.
The derivation of this law follows from:
(i) The assumption that off-equilibrium, and at least for large $\xi$, the equilibrium scaling theory that describes the magnetic response to an external field $H$ holds:
$\begin{split}&M(t,t_{\mathrm{w}};H)=\left[\xi(t+t_{\mathrm{w}})\right]^{-\frac{D}{2}-\frac{\theta}{4}}\\ &\qquad\qquad\times\mathcal{F}\left(H\left[\xi(t+t_{\mathrm{w}})\right]^{-\frac{D}{2}-\frac{\theta}{4}},\frac{\xi(t+t_{\mathrm{w}})}{\xi(t_{\mathrm{w}})}\right).\end{split}$ (56)
The full-aging spin-glass dynamics allows the ratio $\xi(t+t_{\mathrm{w}})/\xi(t_{\mathrm{w}})$ to be considered as approximately constant.
(ii) However, the observation that the Taylor expansion
${\frac{M(t,t_{\mathrm{w}};H)}{H}}=\chi_{1}+\frac{\chi_{3}}{3!}H^{2}+\frac{\chi_{5}}{5!}H^{4}+\mathcal{O}(H^{6})\,,$ (57)
where
$\chi_{2n-1}\propto b_{2n}(T)[\xi(t_{\mathrm{w}})]^{(n-1)D-n\frac{\theta({\bar{x}})}{2}},$ (58)
predicts the paradoxical result that $\chi_{1}\to 0$ when $\xi\to\infty$. In fact, Eq. (58) neglects the contribution of the regular part of the free energy. A better description is
$\chi_{1}={\frac{{\hat{S}}(C_{\text{peak}})}{T}}+{\frac{b_{2}(T)}{\xi^{\theta({\tilde{x}})}/2}},$ (59)
where ${\hat{S}}(C_{\text{peak}})$ is the function appearing in the fluctuation-dissipation relations [89, 204, 122, 36, see Secs. IV.3 and IV.4]. From now on, we shall use the shorthand $\hat{S}$ for ${\hat{S}}(C_{\text{peak}})$.
Figure 36: The non-linear part of the response-time data, $[\ln(t^{\mathrm{eff}}_{H}/t^{\mathrm{eff}}_{H\to 0^{+}})-a_{2}(T)H^{2}]\xi^{\theta(\tilde{x})/2}$, plotted against the scaling variable $[H^{2}\xi^{D-\theta(\tilde{x})/2}]^{2}$, see Eq. (55). On the right, numerical data; on the left the experimental one. See Tables 1-2 for parameters details. Figure taken from Zhai et al. [302].
Figure 36 demonstrates the remarkable collapse of both experimental and numerical data through the scaling law Eq. (55), and solves a thirty-year problem concerning the Zeeman energy. It creates a road map for further comparison between experiments and simulations, such as the investigation of the exotic phenomena of rejuvenation and memory, see Sec. VI.2.
$T_{\mathrm{m}}$ (K) $t_{\mathrm{w}}$ (s) $\xi(t_{\mathrm{w}})/a$ $\theta(\tilde{x})$
Exp. 1 28.50 10 000 320.36 0.337
Exp. 2 28.75 10 000 341.76 0.344
Exp. 3 28.75 20 000 359.18 0.342
Exp. 4 29.00 10 000 391.27 0.349
Table 1: Main parameters for the four experiments reported in Fig. 36. Table taken from Zhai et al. [302].
$T_{\mathrm{m}}$ $t_{\mathrm{w}}$ $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ $t_{\text{max}}$ $\theta(\tilde{x})$ $C_{\mathrm{peak}}(t_{\mathrm{w}})$
Run 1 0.9 $2^{22}$ 8.294(7) $2^{30}$ 0.455 0.533(3)
Run 2 0.9 $2^{26.5}$ 11.72(2) $2^{30}$ 0.436 0.515(2)
Run 3 0.9 $2^{31.25}$ 16.63(5) $2^{32}$ 0.415 0.493(3)
Run 4 1.0 $2^{23.75}$ 11.79(2) $2^{28}$ 0.512 0.422(2)
Run 5 1.0 $2^{27.625}$ 16.56(5) $2^{32}$ 0.498 0.400(1)
Run 6 1.0 $2^{31.75}$ 23.63(14) $2^{34}$ 0.484 0.386(4)
Run 7 0.9 $2^{34}$ 20.34(6) $2^{34}$ 0.401 0.481(3)
Table 2: Parameters for the native numerical simulations reported in Figs. 3639: $T_{\mathrm{m}},~{}t_{\mathrm{w}},~{}\xi(t_{\mathrm{w}})$, the longest simulation time $t_{\text{max}}$, the replicon exponent $\theta$, and the value of $C_{\mathrm{peak}}(t_{\mathrm{w}})$. Table taken from Zhai et al. [302], Paga et al. [232].
The scaling law also provides a vehicle to examine the superposition principle, see Eq. (1) and the discussion in Sec. IV.1. As was briefly explained in the Introduction, this concept was an experimental milestone, describing the connection between the TRM, ZFC, and FC protocols. Exploiting the same techniques developed for the ZFC protocol, Paga et al. [231] focused on the TRM protocol to demonstrate, through the use of correlation lengths, that the superposition principle is valid only in the limit that the magnetic field $H\rightarrow 0^{+}$, see Fig. 37.
Figure 37: Growth of the rescaled magnetization for different experimental protocols: $[M_{\mathrm{ZFC}}(t,t_{\mathrm{w}})+M_{\mathrm{TRM}}(t,t_{\mathrm{w}})]/H$, and $M_{\mathrm{FC}}(0,t_{\mathrm{w}}+t)/H$. The lighter colors refer to the FC case; the darker ones correspond to the quantity ZFC + TRM. The vertical dashed lines indicate the effective times, $t^{\text{eff}}_{H}$ , associated with each case. The violation of Eq. (1) is evident. Figure taken from Paga et al. [231].
The argument is as follows. In the out-of-equilibrium regime, we can connect the growth of $\xi_{\mathrm{micro}}$ to time as in Eq. (49). Additionally, we can relate $\xi_{\mathrm{micro}}$ to an external magnetic field:
[…]
where $D$ is the spatial dimension and $\theta$ is the replicon.
Whether in the presence (TRM) or absence (ZFC) of an external magnetic field, the study begins with the calculation of $\xi_{\mathrm{micro}}$ using Eq. (24). The presence of an external magnetic field slows down the growth of $\xi_{\mathrm{micro}}$. Because $\xi_{\mathrm{micro}}^{\mathrm{ZFC}}$ and $\xi_{\mathrm{micro}}^{\mathrm{TRM}}$ approach the $H^{2}\to 0^{+}$ limit with a linear slope, their difference can be expressed as
$1-\frac{\xi_{\mathrm{micro}}^{\mathrm{TRM}}(t_{\mathrm{w}},H)}{\xi_{\mathrm{micro}}^{\mathrm{ZFC}}(t_{\mathrm{w}})}=A(t_{\mathrm{w}},T)[\xi_{\mathrm{micro}}^{\mathrm{ZFC}}(t_{\mathrm{w}})]^{D-\theta/2}H^{2}\,.$ (61)
A scaling behavior emerges, see Fig. 38.
Figure 38: Scaling relation according to Eq. (61). The main plot is in log-log scale; the insert has a linear scale for the ordinate. Figure taken from Paga et al. [231].
The same result can be obtained following a different strategy. Because of Eq. (55), the decay of the effective time $t^{\text{eff}}_{H}$ can be connected to $\xi_{\mathrm{micro}}$. Thus, the difference in the decay of $t^{\text{eff}}_{H}$ between the TRM and ZFC protocols can be written as
$\begin{split}\frac{\ln t^{\text{eff,ZFC}}_{H}-\ln t^{\text{eff,TRM}}_{H}}{\xi_{\mathrm{micro}}(t_{\mathrm{w}})^{D-\theta/2}}=&-[K^{\mathrm{ZFC}}\!-\!K^{\mathrm{TRM}}]H^{2}\\ &+\mathcal{O}(H^{4})\,.\end{split}$ (62)
The results in Fig. 39 show that scaling behavior emerges. By combining this with Eq. (62), the failure of the superposition principle, Eq. (1), in finite magnetic fields is shown to be a result of the lowering of the free-energy barrier in the presence of an external magnetic field.
Figure 39: The rescaled quantity $T(\delta\ln t^{\text{eff}}_{H})/[\xi_{\mathrm{micro}}(t_{\mathrm{w}})]^{D-\theta/2}$ as a function of $H^{2}$. Figure taken from Paga et al. [231].
V.4 Properties in the vicinity of the critical temperature
A scaling law for the magnetic field valid for large fields and close to $T_{\mathrm{g}}$ was developed in Zhai et al. [300]. It was based on the assumption that a scaling theory that describes the magnetic response to an external field $H$ in equilibrium [238, 14] also holds in the non-equilibrium regime. Defining an “effective” response time, $t^{\text{eff}}_{H}$ in the same sense as Eq. (17) above, Eqs. (5558) can be written as
$\ln\,t_{H}^{\text{eff}}=a_{0}+a_{2}H^{2}+a_{4}H^{4}+a_{6}H^{6}+{\mathcal{O}}(H^{8}),$ (63)
where
$a_{n}\propto b_{2n}(T)\xi^{|nD-(n+1)\theta(\bar{x})|/2}.$ (64)
The terms in Eq.  (64) are defined in the text following Eq. (55). Fits of Eqs. (57) and (58) to the data exhibit non-linearities in $H^{2}$, both from experiments and simulations, and are exhibited in Zhai et al. [300].
Figure 40: Effective time $t_{\mathrm{w}}^{\text{eff}}$ extracted from the time at which $S(t)$ peaks vs temperature for waiting times $t_{\mathrm{w}}$ from 300 to 100 000 s for a 6 at.CuMn single-crystal sample ($T_{\mathrm{g}}=31.5$ K). Reproduced from Kenning [171]
A spectacular example of the need for the scaling law is the collapse of $t^{\text{eff}}_{H}$ near $T_{\mathrm{g}}$ [171]. Fig. 40 exhibits his extraction of $t^{\text{eff}}_{H}$ for a CuMn 6 at.($T_{\mathrm{g}}=31.5$ K) as $T\rightarrow T_{\mathrm{g}}$ from below.
This collapse is evidence of the importance of the higher-order terms in the scaling law [Eqs. (57) and (58)]. The values of $a_{2}$ and $a_{4}$ were found by Zhai et al. [302] to be of opposite sign, so that the $H^{4}$ term is of opposite sign to the $H^{2}$ term in the expansion of Eq. (63). The power of the coherence length in the $H^{4}$ term is roughly twice the power of the coherence length in the $H^{2}$ term. Hence, as one approaches $T_{\mathrm{g}}$ [and $\xi_{\mathrm{micro}}$ increases, 34], the $H^{4}$ term rapidly grows compared to the $H^{2}$ term, sharply diminishing $t^{\text{eff}}_{H}$.
A quantitative fit is underway at the time of this review [171]. For the moment, we believe the scaling law exhibited in Eqs. (63) and (64) is sufficient to explain the collapse of $t^{\text{eff}}_{H}$ as $T\rightarrow T_{\mathrm{g}}$ from below.
V.5 Brief summary of dynamics in a field
The equilibrium and out-of-equilibrium dynamics of the 3D Ising spin glass in the presence of an external magnetic field was addressed in Baity-Jesi et al. [25] using numerical simulations with the Janus I dedicated supercomputer. The next paragraphs briefly review the main findings of this study.
The dynamics appear to be activated, in agreement with the droplet-model predictions that no phase transition occurs in a field. However the detailed dependence of the observables with time differs from that predicted by the droplet model. In particular, the data follows a super-Arrhenius law at temperatures as low as $T=0.36T_{\mathrm{g}}^{(H=0)}$. The large value of the correlation length at low temperatures indicates that the lower critical dimension is at most slightly greater than three.
Furthermore, equilibrium and out-of-equilibrium approaches allow the determination of a kind of transition line. However, the correlation length does not diverge on this line, against replica symmetry breaking (RSB) predictions. In addition, if the transition is driven by a $T=0$ fixed point (as in the droplet model) one should expect some kind of activated dynamics also in RSB. Thus, in this scenario, the identified transition line could be interpreted as a dynamic glass transition with the correlation length diverging at lower temperatures.
Finally, Baity-Jesi et al. [25] discussed a scenario motivated by the study of supercooled liquids. On a qualitative level, the dynamic transition could be interpreted as a mode-coupling temperature, and, at lower temperatures, a crossover to activated dynamics.
In summary, the dynamical behavior in and out of equilibrium of the three-dimensional spin glass in a magnetic field exhibits large values of the correlation length. But this behavior is not compatible with the onset of a continuous phase transition (a de Almeida Thouless line). These facts, together with the results of equilibrium numerical simulations [26, 11], suggest that the lower critical dimension of the Ising spin-glass model in a field is greater than three but less than four [see also 232].
VI Off-equilibrium dynamics: non-linear response when varying the temperature
Rejuvenation and memory experiments [161] are among the most profound in the field of spin glasses. These experiments suggested that the study of magnetic responses in temperature-varying protocols could demonstrate some of the most interesting features of a complex free-energy landscape (most notably, the multiplicity of equilibrium states and temperature chaos). Once again, it is necessary to relate equilibrium landscapes to non-equilibrium experiments. This section reports on a continuing effort, originating in Bouchaud et al. [67] and still ongoing [31, 125, 230], to learn about the geometry of the free-energy landscape from temperature-varying experiments.
This section does not take a chronological view. Rather, it intertwines experimental and theoretical progress on different aspects of the problem. Sec. VI.3 is somewhat different, because it deals with a memory of a different kind (the Mpemba effect), which has been theoretically predicted but not yet seen experimentally in spin glasses.
VI.1 Temperature chaos in the off-equilibrium context
Temperature chaos was introduced as an equilibrium phenomenon in Secs. II.4 and III, while experiments are always conducted out of equilibrium. This section will explore how one can translate the concept of temperature chaos to the experimental setting and, then, how it can be investigated in non-equilibrium simulations.
VI.1.1 The experimental point of view
Figure 41: The reference and the temperature cycling curves for $M_{\text{FC}}(t)$ of a CuMn 6 at.sample for $T_{1}=18$ K, and $t_{\mathrm{w1}}=10^{4}$ s. The temperature cycling curve is shifted by $\delta t$ to overlap the reference curve. In the reversible range, (a) and (b), the cycling curve can be overlapped with the reference curve over the whole period of measurement, $\approx 7\times 10^{4}$ s. In the chaotic range, (c) and (d), the cycling curve can only be partially overlapped. Hence, temperature chaos sets in for $T_{1}=18$ K, $t_{\mathrm{w1}}=10^{4}$ s at $\Delta T\geq 450$ mK. Reproduced from Fig. 1 of Zhai et al. [301].
Bray and Moore [70] introduced a length scale for temperature chaos upon a temperature change $\Delta T$:
$\ell_{\text{c}}=a_{0}\bigg{[}{\frac{T_{1}}{|\Delta T|}}\bigg{]}^{1/(D_{\text{s}}/2-y)}$ (65)
where $a_{0}$ is the average distance between magnetic ions, $\Delta T=T_{1}-T_{2}$ is the temperature change, $T_{1}$ and $T_{2}$ being the initial and final temperatures, respectively, and $(D_{\text{s}}/2-y)$ is the exponent combination for temperature chaos (Sec. II.4). Thus, at a given coherence length $\xi(t_{\mathrm{w}},T)$, the condition for the onset of chaos is
$\ell_{\text{c}}=\xi(t_{\mathrm{w}},T),$ (66)
such that, for $\ell_{\text{c}}\geq\xi(t_{\mathrm{w}},T)$ the system is reversible, but for $\ell_{\text{c}}\leq\xi(t_{\mathrm{w}},T)$, the system has entered a chaotic state.
The search for temperature chaos in spin glasses has had a checkered history. The first systematic report [163] utilized both positive and negative temperature shifts for $\Delta T$. Arai et al. [18] were able to change bond strengths to access the chaotic state. Jönsson et al. [163] found for small $\Delta T$ that aging was “accumulative” but for larger $\Delta T$ a transition to “non-accumulative” aging was observed.
We associate accumulative to reversible behavior, and non-accumulative to chaos. Jönsson et al. [163] used ZFC measurements to extract the temperature ranges for accumulative and non-accumulative aging. They estimated the chaos exponent $(D_{\text{s}}/2-y)$ as
$1/(D_{\text{s}}/2-y)_{\text{J}}=2.65\pm 0.5$ (67)
where the subscript J refers to their value. Unfortunately, the accepted value for $1/(D_{\text{s}}/2-y)$ is near unity (recall Sec. II.4). The origin of their error has been speculated [301] to arise from the presence of chaos induced by the change of magnetic field in a ZFC protocol. It is known that magnetic-field change can induce chaos [181, 56], so that it may have interfered with their extraction of the chaos exponent.
The problem, therefore, is how to extract $1/(D_{\text{s}}/2-y)$ from experiment without changing a magnetic field. A solution was developed by Zhai et al. [301], who utilized the (small but finite) time dependence of the field-cooled magnetization, $M_{\text{FC}}(t)$. In this protocol, the magnetic field remains constant, so that only a change in temperature can contribute to chaos.
The protocol was first to establish a reference curve. A CuMn 6 at.single crystal ($T_{\mathrm{g}}=31.6$ K) was cooled from 40 K to the measurement temperature $T_{1}=18$ K in a constant field $H=40$ Oe at 10 K/min. After the temperature was stabilized, the reference magnetization $M_{\text{FC}}^{\text{ref}}(t,T_{1})$ was recorded. In the temperature-cycling protocol, the sample was cooled from 40 K to $T_{1}=18$ K at 10 K/min, where the magnetization was recorded for a duration of $10^{4}$ s. The temperature was then lowered to $T_{2}=T_{1}-\Delta T$. After aging for $10^{3}$ s, the temperature was raised back to $T_{1}=18$ K, and the magnetization $M_{\text{FC}}^{\text{cyclic}}(t)$ recorded.
If there is no chaos, the two time-dependent magnetizations, $M_{\text{FC}}^{\text{ref}}(t,T_{1})$ and $M_{\text{FC}}^{\text{cyclic}}(t)$, can be overlapped directly by shifting $M_{\text{FC}}^{\text{cyclic}}(t)$ by $\delta t$ [see Fig. 41, curves (a) and (b)], and aging can be taken as cumulative and reversible. However, if chaos is present, $M_{\text{FC}}^{\text{cyclic}}(t)$ cannot be made to overlap with $M_{\text{FC}}^{\text{ref}}(t,T_{1})$ by shifting $M_{\text{FC}}^{\text{cyclic}}(t)$ by $\delta t$ [see Fig. 41, curves (c) and (d)].
Figure 41 displays the results of these experiments. Comparing Fig. 41 (b) with Fig. 41 (c) shows that temperature chaos begins between 17.55K and 17.54 K, or between $\Delta T=0.45$ K and 0.46 K, for an unprecedented accuracy of 10 mK. The amplitude of the lack of overlap increases with increasing $\Delta T$, as can be seen in Fig. 41 (d) where $\Delta T=12$ K. Zhai et al. [301] display additional values for $\Delta T$ for $T_{1}=18$ K, and for different $T_{1}=17-14$ K. Having these different $T_{1}$, and the concomitantly different $\Delta T$ for the onset of temperature chaos, enables the extraction of $1/(D_{\text{s}}/2-y)\approx 1.1$, well in the range of accepted values for $(D_{\text{s}}/2-y)$ from theoretical calculations and simulations for the 3D Ising model [see 301, for a list of these values].
VI.1.2 The numerical point of view
The crucial elements to successfully observe temperature chaos in non-equilibrium simulations are three: the capacity to reach long times (enabled by specialized hardware, see Sec. II.7), the idea that temperature chaos is a rare event [111, see Sec. III], and the statics-dynamics equivalence [41, 43, 13, 36, see Secs. IV.3 and IV.4].
Baity-Jesi et al. [30] performed numerical simulations of the EA model in three dimensions with a lattice size of $L=160$. The numerical protocol consisted of a direct quench from a random configuration, effectively at infinite temperature, to a working temperature $T_{\mathrm{m}}<T_{\mathrm{g}}$, where the system is left to relax for a time $t_{\mathrm{w}}$.
Figure 42: Randomly chosen spheres of a sample of size $L=160$ are depicted with a color code depending on $1-X$ ($X$ is the chaotic parameter computed for spheres of radius $r=12$ and $\xi_{\mathrm{micro}}=12$ and temperature $T_{1}=0.7$ and $T_{2}=1.0$). For visualization purposes, spheres are represented with a radius $12(1-X)$, so that only fully chaotic spheres (i.e., $X=0$) have their real size. Figure from [30].
At this point, it is necessary to, somehow, compare configurations with different temperatures $T$ of the thermal bath. The natural choice is to use the chaotic parameter $X^{J}_{T_{1},T_{2}}$, see Eq. (13). While at equilibrium computing the chaotic parameter is straightforward, in off-equilibrium dynamics there is an extra parameter to play with: it is necessary to select the time $t$ of the compared configurations. The statics-dynamics equivalence suggests that, if in equilibrium the relevant length scale is the linear size $L$, the non-equilibrium system should be governed by the coherence length $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$, whose growth strongly depends on $T$ (see Sec. IV.2). Thus, the authors used the coherence length of the system as a proxy of its aging state and compared configurations at different times by keeping $\xi_{T_{1}}(t_{\mathrm{w1}})\approx\xi_{T_{2}}(t_{\mathrm{w2}})$. This strong condition will be discussed again in Sec. VI.2.2.
When studying the chaotic parameter as a global observable, the authors observed almost no signal. Again, the statics-dynamics equivalence suggests that the non-equilibrium system with $L\gg\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ can be considered as a collection of equilibrated systems with $L\sim\xi_{\mathrm{micro}}(t_{\mathrm{w}})$. Therefore, computing global observables implies averaging over equilibrated regions. As explained in Sec. III.1, this averaging washes the chaotic signal away.
Hence, the study of temperature chaos has to be local. The authors defined spherical regions of size $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ and computed within them a chaotic parameter
$X^{s,r}_{T_{1},T_{2}}(\xi)=\frac{\bigl{\langle}\bigl{[}q^{s,r}_{T_{1},T_{2}}\bigr{]}^{2}\bigr{\rangle}_{J}}{\bigl{(}\bigl{\langle}\bigl{[}q^{s,r}_{T_{1},T_{1}}\bigr{]}^{2}\bigr{\rangle}_{J}\bigl{\langle}\bigl{[}q^{s,r}_{T_{2},T_{2}}\bigr{]}^{2}\bigr{\rangle}_{J}\bigr{)}^{1/2}}\,,$ (68)
where the superindex $\{s,r\}$ labels the specific sphere $s$ and the radius of the considered sphere $r$ for which we are computing the chaotic parameter and the thermal averages $\langle\cdots\rangle_{J}$ are as in Eq. (13). This local observable emphasizes the heterogeneity of the phenomenon, as shown in Fig. 42, in which there is a vast majority of non-chaotic spheres (blue-like colors) and specific regions with very chaotic spheres (red-like colors).
Figure 43: Distribution function $F(X,T_{1},T_{2},\xi_{\mathrm{micro}},r)$ for $T_{1}=0.625$ and $T_{2}=0.9$, for spheres of fixed radius $r=4$ (top) and fixed coherence length $\xi_{\mathrm{micro}}=9$ (bottom). Data taken from Baity-Jesi et al. [30].
A more quantitative study was conducted in Baity-Jesi et al. [30] by computing the probability distribution function of the chaotic parameter of the spheres,
$F(X,T_{1},T_{2},\xi,r)=\mathrm{Prob}\bigl{[}X^{s,r}_{T_{1},T_{2}}(\xi)<X\bigr{]}.$ (69)
Three main results can be extracted from this computation. First, the vast majority of the spheres exhibit no chaos. See, for example, Fig. 43 (top), $\xi_{\mathrm{micro}}=9$ curve, approximately $90\%$ of the spheres have $X^{s,r}_{T_{1},T_{2}}>0.9$. Next, the chaotic signal monotonically increases with the coherence length $\xi_{\mathrm{micro}}$. Last, for each coherence length, there exists a value for the radius of the sphere that maximizes the chaotic signal [see Fig. 43 (bottom)].
This analysis, though very rich and interesting, is also afflicted by an excess of information: there are a lot of variables that can be explored. To clarify the situation and focus on the strong chaotic signal, the authors fixed the probability level $F$ and the coherence length $\xi$, studying the change of the chaotic signal with the size of the sphere. Full details of the procedure can be found in Baity-Jesi et al. [30] and Moreno-Gordo [215].
As sketched above, it was found that the chaotic signal peaks for a specific value of the radius of the studied sphere. The analysis of the peaks found two main results. First, the optimum radius to study temperature chaos is proportional to the coherence length $\xi_{\mathrm{micro}}$. Second, there exists a characteristic coherence length $\xi_{\mathrm{micro}}^{*}$ that points to a crossover between a weak-chaotic regime and a strong-chaotic regime. Moreover, since $\xi_{\mathrm{micro}}^{*}$ indicates the crossover between weak and strong chaos, it must be the non-equilibrium analog of the equilibrium chaotic length $\ell_{\text{c}}$ (see Sec. II.4). The authors found that the crossover length $\xi_{\mathrm{micro}}^{*}$ admits a scaling with the temperature difference $T_{2}-T_{1}$ of the form
$\xi_{\mathrm{micro}}^{*}(T_{1},T_{2},F)\propto(T_{2}-T_{1})^{-1/(D_{\text{s}}/2-y)_{\mathrm{NE}}}\,$ (70)
very similar to the scaling of the chaotic length $\ell_{\text{c}}$. Fernandez et al. [111] found a value of the exponent $(D_{\text{s}}/2-y)=1.07(5)$ at equilibrium while Baity-Jesi et al. [30] found $(D_{\text{s}}/2-y)_{\mathrm{NE}}=1.19(2)$ out of equilibrium. This agreement supports the idea of a crossover length.
The crossover length $\xi^{*}$ turned out to be fundamental to understanding temperature chaos out of equilibrium and allows the estimation of the necessary conditions to observe effects that deeply depend on the presence of a strong chaotic regime, such as memory and rejuvenation.
In summary, temperature chaos is real, at least from an experimental point of view. The onset is clear from Fig. 41, and the extraction of $(D_{\text{s}}/2-y)$ gives a value consistent with theory and simulations. What remains is the nature of temperature chaos: is it a rapid rise from onset, or does it smoothly increase across a broad range of $\Delta T$? Another unanswered question is its magnitude. Does the chaotic state fully replace the reversible or cumulative aging state, or is it small compared to it? The next section, on memory, will provide at least a partial answer.
VI.2 Rejuvenation and memory
VI.2.1 An experimental approach
One of the remarkable features of spin-glass dynamics is their behavior under aging and temperature cycling, namely rejuvenation and memory. These properties were first exhibited by Jonason et al. [162] and are reproduced in Fig. 44. The imaginary part of the ac magnetic susceptibility, $\chi^{\prime\prime}(\omega)$, of the insulating spin glass CdCu1.7In0.3S4 ($T_{\mathrm{g}}=16.7$ K) as a function of temperature $T$ was first measured without aging, providing the reference curve (solid line in Fig. 44). Then the experiment was repeated, but this time as the temperature was lowered, it was stopped at $T=12$ K, where the system aged for 7 h. Then the temperature was lowered to the final $T=5$ K. Remarkably, $\chi^{\prime\prime}(\omega)$ returned to the reference curve after a small change in $T$. This behavior was termed “rejuvenation”: The measured $\chi^{\prime\prime}(\omega)$ behaved as if it had never been aged at $T=12$ K! This effect has been attributed to temperature chaos, but there remains need for a definitive proof of this conjecture.
Figure 44: The out-of-phase susceptibility $\chi^{\prime\prime}$ of the CdCr1.7In0.3S4 spin glass. The solid line is measured upon heating the sample at a constant rate of 0.1 K/min (reference curve). Open diamonds: the measurement is done during cooling at this same rate, except that the cooling procedure has been stopped at 12 K during 7 h to allow for aging. Cooling then resumes down to 5 K: $\chi^{\prime\prime}$ is not influenced and goes back to the reference curve (chaos). Solid circles: after this cooling procedure, the data is taken while reheating at the previous constant rate, exhibiting memory of the aging stage at 12 K. Reproduced from Fig. 1 of Jonason et al. [161].
What was seen next was even more unexpected. As the system was heated from its lowest temperature, $\chi^{\prime\prime}(\omega)$ retraced its behavior, reproducing the drop seen during aging when the temperature was decreasing. This was termed “memory”. That it occurred after the system had rejuvenated makes it remarkable. For if rejuvenation is (and was) ascribed to a transition to a chaotic state, how could it “remember” what has occurred at a previous higher temperature after aging?
Figure 45: Measurements of the ac susceptibility displaying memory as a function of temperature. The imaginary part of the 1 Hz magnetic susceptibility, $\chi^{\prime\prime}(\omega)$. The dashed data are the reference curve measured while continuously cooling the sample at 1 K/min. the dot-dashed data include waiting at $T_{2}=22.5$ K for a time $t_{\mathrm{w1}}=1$ h and rejuvenation upon lowering the temperature. The temperature cooling rate is also 1 K/min. The solid curve is the heating data taken at a rate of 1 K/min that exhibits the memory effect at $T_{1}$. The inset shows an example of a double-waiting-time experiment. Reproduced from Fig. 1 of Freedberg et al. [125].
The breakthrough in our understanding of the coexistence of rejuvenation and memory originated with the experiments of Freedberg et al. [125] and the simulations of the Janus Collaboration [31, 230]. The former quantified memory through
${\mathcal{M}}={\Delta\chi_{H}^{\prime\prime}}/{\Delta\chi_{C}^{\prime\prime}}.$ (71)
Using this definition for the amount of memory, see Fig. 45, Freedberg et al. [125] presented a quantitative analysis of the effects of two wait times $t_{\mathrm{w1}}$ and $t_{\mathrm{w2}}$ at various temperatures $T_{1}$ and lower temperatures $T_{2}$ based on a real-space model for the growth of the correlations at the two temperatures.
The outstanding element of this work was the development of a phenomenological model of memory with three adjustable parameters: a scaling constant, a length scale, and an exponent. The scaling constant was of order unity and the exponent, though not fully understood, was reasonable. After developing the physics of the model below, more will be said regarding the length scale.
Figure 46: ${\mathcal{M}}$ vs waiting time with four different first waiting temperatures. For both (a) and (b), the first and second waiting temperatures are the same: $T_{1}=18,~{}22.5,~{}27.5$, and 30 K and $T_{1}-T_{2}=$ 4 K. In both cases either the first (a) or second (b) waiting times were varied from 0 - 6 h (the horizontal axis), while the other waiting time was fixed at 3 and 1 h, respectively. Closed (open) markers indicate a variation of the first (second) waiting time. Reproduced from Fig. 2 of Freedberg et al. [125].
Figure 47: Memory vs waiting time data plotted against a model for memory loss predicted by [125]. Additionally, the three-starred data points use coherence lengths extracted from dc experiments with the ac values of memory. The data collapse supports the model. $T_{1}$ (18, 22.5, 27.5 and 30 K) and $T_{1}-T_{2}=4$ K. In both cases either the first (closed markers) or second (open markers) waiting times are varied from 0 to 6 h, while the other waiting time is fixed at 3 and 1 h, respectively. Statistical error bars are present on each data point, though they are typically smaller than the marker. Reproduced from Fig. 3 of Freedberg et al. [125].
The model quantitatively describes how the correlations at the first waiting temperature grow, resulting in a decreasing susceptibility. Quenching to the second wait temperature, those correlations remain but new correlations at the second temperature begin to grow and “eat” into the higher-temperature correlations. Upon returning to the first waiting temperature, the original correlations are present but now reduced in magnitude. The remarkable agreement of this model with the data is illustrated in Fig. 46, where it is important to note that the same adjustable parameters are used for all the exhibited data.
As to the length scale, their data show there must be another length scale at play other than the coherence length that is on the order of 100 nm. The origin of this length scale is not understand but may be related to the equilibrium correlation length.
All the data is considered at once in Fig. 47, which shows how the memory scales with $\xi_{1}(\xi_{2}/\xi_{1})^{3}$, where 1 and 2 refer to the higher and lower temperatures, respectively. This remarkable data collapse shows the memory depends explicitly on the coherence length at both temperatures. It also shows the importance of another length scale in the problem. The data in this figure are from 52 double-waiting-time experiments.
This work shows a physically reasonable model that explains memory and the interference between two different waiting-time temperatures. Complementary data were reported by Paga et al. [230] and by He and Orbach [146] “through the ‘prism’ of the coherence length.”
Figure 48: The abbreviations N, R, and C stand for native, rejuvenation, and cycle, respectively. In N3, the temperature is lowered from above $T_{\mathrm{g}}=41.6$ K, for a CuMn single crystal 8 at.%, to the lower temperature $T_{2}$ (here, $T_{2}=26$ K, $t_{\mathrm{w2}}=3$ h), and the effective waiting time, given by the peak in $S(t)$ defined by Eq. (5), is measured for different magnetic fields. Next, the system is cooled from above $T_{\mathrm{g}}$ to $T_{1}=30$ K and aged for 1 h. The temperature is then dropped to $T_{2}=26$ K, and aged for 3 h. The points are labeled $R_{1}$. As can be seen, the two procedures yield nearly exactly the same $t^{\mathrm{eff}}_{\mathrm{w}}$ for all values of $H^{2}$, independent of the aging at $T_{1}$. This is a clear demonstration of rejuvenation. Memory is measured through a full temperature cycle, from $T_{1},t_{\mathrm{w1}}\rightarrow T_{2},t_{\mathrm{w2}}\rightarrow T_{1}$ where $t^{\mathrm{eff}}_{\mathrm{w}}$ is measured. The text discusses the physical meaning for the three protocols $C_{1},~{}C_{2}$ and $C_{3}$. See Tab. 3 for parameter details. Reproduced from Fig. 2 of Paga et al. [230].
Protocol $T_{1}$ (K) $t_{\mathrm{w1}}$ (h) $T_{2}$ (K) $t_{\mathrm{w2}}$ (h) $T_{\mathrm{m}}$ (K) $\xi/a$
N1 30 1 30 13.075
N2 26 1/6 26 8.1011
N3 26 3 26 11.961
N4 16 3 16 6.3271
R1 30 1 26 3 26 11.787
C1 30 1 26 1/6 30 12.621
C2 30 1 26 3 30 12.235
C3 30 1 16 3 30 12.846
Table 3: Parameters of the experiments shown in Fig. 48. Table taken from Paga et al. [230].
Upon cooling the spin glass from above $T_{\mathrm{g}}$ to the first measuring temperature $T_{1}$, the system ages for a time $t_{\mathrm{w1}}$. As a consequence, the spin-glass coherence length grows from nucleation to $\xi(t_{\mathrm{w1}},T_{1})$. When the temperature is then lowered to $T_{2}$, the correlations created at $T_{1}$ are frozen. This concept was first introduced by Bouchaud et al. [67]. In accord with Freedberg et al. [125], when the system is aged at $T_{2}$ for a time $t_{\mathrm{w2}}$, the system is now in a chaotic state that has nothing to do with the state created at $T_{1},t_{\mathrm{w1}}$. Its coherence length encompasses a coherent volume that interferes with the coherent volume created at $T_{1},t_{\mathrm{w1}}$. Hence, when heating back to $T_{1}$, the net correlated volume is the difference between the two correlated volumes.
This is exhibited in the lower part of Fig. 48. The $C_{n}$ are separate temperature and waiting-time cycles, and illustrate unequivocally the relationship between memory and competing correlation volumes. Each of the $C_{n}$ has three steps: 1) the system is “prepared” at $T_{1}=30$ K by waiting for the same time $t_{\mathrm{w1}}=1$ h. 2) The temperature is dropped to $T_{2}$ and the system is aged for $t_{\mathrm{w2}}$. 3) The system is heated back to $T_{1}=30$ K and $t^{\mathrm{eff}}_{\mathrm{w}}$ measured. $C_{1}$ sets $T_{2}=26$ K and $t_{\mathrm{w2}}=1/6$ h. $C_{2}$ sets $T_{2}=26$ K and $t_{\mathrm{w2}}=3$ h. $C_{3}$ sets $T_{2}=16$ K and $t_{\mathrm{w2}}=3$ h.
Memory is quantified by comparing the magnitude of the correlation volume measured at step 3) with that of the native protocol [230]:
${\cal C}_{\text{Zeeman}}={N_{\text{s}}^{\text{cycle}}}/{N_{\text{s}}^{\text{native}}}\,,$ (72)
where $N_{\text{s}}$ is the number of correlated spins, as in Eq. (11). If the two volumes are the same, memory is perfect. If, after step 3), the measured correlation volume is smaller than the initially prepared volume at step 1), memory is lessened, a direct result of the interference between the two states. The slopes in Fig. 48 are steeper the larger the coherent volume [Eqs. (10) and (11)].
Considering $C_{1}$, the system has transitioned from the prepared state at $T_{1}$ into a chaotic state at $T_{2}$, but only for a short time (1/6 h). The coherent volume in the chaotic state has grown during this time, so that its interference with the initially prepared coherent volume is significant. Hence, the memory is diminished, as exhibited by the shallower slope as compared to the native slope exhibited in Fig. 48.
Now, $C_{2}$ increases $t_{\mathrm{w2}}$ to 3 h, so that the coherent volume in the chaotic state can grow beyond its value in $C_{1}$. This should lead to a larger interference, hence smaller memory, leading to a more shallow slope than found for $C_{1}$. Its behavior is seen in Fig. 48.
Finally, the $C_{3}$ protocol has the same $t_{\mathrm{w2}}$ as $C_{2}$, but the temperature drop to $T_{2}$ is much greater. At such a low temperature, the growth of the coherent volume of the chaotic state is very slow (almost none in this experiment), so there should be almost no interference with the coherent volume created at $T_{1}$. This is exhibited in Fig. 48, where the slope of $C_{3}$ is close to the native slope.
Figure 49: The ZFC numerical experiment measuring rejuvenation and memory. At time $t_{\mathrm{w}}$, a magnetic field $H=0.01$ is applied and the magnetic density, $M_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$, is recorded. Left panels show aging. The relaxation function $S_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$ is displayed for the native runs at the warmer, $T_{1}=0.9$, and colder, $T_{2}=0.5$, temperatures . In the temperature-varying protocol (schematized by the green arrows), after a waiting time $t_{\mathrm{w1}}=2^{31.25}$ at the cold temperature $T_{1}=0.9$, the system is quenched to the cold one $T_{2}=0.5$ Then, the system relaxes at $T_{2}$ for an additional time, after which the magnetic field is switched on and the function $S_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$ is measured (bottom-center). Finally, after the waiting time $t_{\mathrm{w}}^{\uparrow}=t_{\mathrm{w1}}+t_{\mathrm{w2}}=2^{32.25}$ (i.e., the system has spent half its life at the initial temperature $T_{1}$ and half at the colder temperature $T_{2}$ without a field), the spin glass is suddenly heated back to $T_{1}$. The system is allowed to relax for a short time, $t_{\mathrm{w}}=2^{10}\ll t_{\mathrm{w1}}$, after which the magnetic field is switched on. The $S_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$ measured after the jump back (top-right) has a peak very similar to the one before the first jump (top-left), evincing the memory of the aging at the initial temperature $T_{1}$, notwithstanding the rejuvenation observed when staying at the lower temperature $T_{2}$. Figure taken from Baity-Jesi et al. [31].
Figure 50: Aging dynamics is controlled by three length scales. The solid lines and filled circles are for $\xi_{\mathrm{Zeeman}}(t_{\mathrm{w}},T)$, the dashed lines and empty circles are for $\xi_{\mathrm{micro}}(t_{\mathrm{w}},T)$, and the dotted lines and empty squares are for $\zeta(t_{1},t_{2})$. Left: for the native protocols and jump runs at $T_{2}=0.5$, $\xi_{\mathrm{Zeeman}}(t_{\mathrm{w}})$ follows quite closely the behavior of $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$. The jump to $T_{2}=0.5$ satisfies the chaos requirement in Eq. (73) and hence exhibits rejuvenation. When the system jumps back to $T_{1}=0.9$ (i.e., $T_{1}=0.9\to T_{2}=0.5\to T_{1}=0.9$), $\xi_{\mathrm{Zeeman}}^{\mathrm{jump-back}}(t_{\mathrm{w}})$ goes back to its original value $\xi_{\mathrm{micro}}^{\mathrm{native}}(t_{\mathrm{w}})$ after an extremely short time (memory). Instead, $\xi_{\mathrm{Zeeman}}(t_{\mathrm{w}})$ never becomes small for the jump protocols with $T_{2}=0.7$. The $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ for the jump runs are superimposed (the spins are frozen for this length scale). Right: the size of the regions undergoing coherent rearrangements when evolving from the initial to the final time, $\zeta(t_{1},t_{2})$, is much smaller than $\xi_{\mathrm{micro}}(t_{\mathrm{w}})$ for all jump protocols. The two times at which the correlation length, $\zeta(t_{1},t_{2})$, is evaluated are $t_{1}=t_{\mathrm{w}}$ and $t_{2}=2t_{\mathrm{w}}$. Inset: comparison of $\zeta(t_{1}=t_{\mathrm{w}},t_{2}=2t_{\mathrm{w}})$ and $\xi_{\mathrm{Zeeman}}(t_{\mathrm{w}})$. For native runs, the two lengths are approximately equal. Instead, for the jump protocols $1.8\,\zeta(t_{1}=t_{\mathrm{w}},t_{2}=2t_{\mathrm{w}})$ is shown. Indeed, for $T_{2}=0.5$ only, and using an appropriate scaling factor of approximately 2, it is clear that $\zeta^{\mathrm{jump}}(t_{1}=t_{\mathrm{w}},t_{2}=2t_{\mathrm{w}})$ can be made to coincide with $\xi_{\mathrm{Zeeman}}^{\text{jump}}$. Figure taken from Baity-Jesi et al. [31].
These three temperature cycles are strong evidence for this interpretation of memory. This is at odds with Bouchaud et al. [67], who argue that the coherent volume returns to the native value upon reheating. Adjusting $t_{\mathrm{w2}}$, one can change the value of the coherent volume when returning to $T_{1}$ at will, from memory loss to complete loss of memory. This behavior is born out in Fig. 48, and also in the independent experiments and model of Freedberg et al. [125].
VI.2.2 Rejuvenation and memory in computer experiments
This section presents the full numerical evidence for the phenomena of rejuvenation and memory, which builds on all the physical insight derived from the results in the previous sections of this review, most notably the scaling law in Sec. V.3 and the formalization of temperature chaos out of equilibrium (see Sec. VI.1.2).
Inspired by experiments on single-crystal CuMn, the Janus Collaboration [31] conducted extensive numerical simulations to mimic ZFC protocols with temperature cycles, as in Fig. 44. They implemented a fixed-temperature or native protocol as well the jump or jump-back temperature-varying protocols. The native runs were already described in Sec. V, the other two are as follows:
  • Jump run: The starting random spin configuration is instantaneously placed at the hot temperature $T_{1}<T_{\mathrm{g}}$ and allowed to relax for time $t_{\mathrm{w1}}$ without a field, achieving a large $\xi_{\mathrm{micro}}(t_{\mathrm{w1}})\sim 17\,\mathrm{a}_{0}$. Then, the temperature abruptly drops to the cold temperature $T_{2}$. After an additional time $t_{\mathrm{w2}}$, the magnetic field is switched on and the relaxation function $S_{\mathrm{ZFC}}$, Eq. (2), is measured.
  • Jump-back run: After the system relaxes at the cold temperature $T_{2}$, it returns to the hot temperature $T_{1}$ before turning on the magnetic field.
Recall that in native runs the peak of $S_{\mathrm{ZFC}}(t,t_{\mathrm{w}};H)$ defines $t^{\mathrm{eff}}_{\mathrm{w}}\simeq t_{\mathrm{w}}$ [compare Fig. 35 with Fig. 49 (left)].
The experimental results of Li et al. [191] suggest that the temperature chaos in a CuMn sample is weak whenever $T_{1}-T_{2}$ is small. It follows that in a simulation, chaotic spheres will be too rare to affect the overall sample relaxation substantially unless
$\left.\frac{T_{1}}{T_{2}}\right|_{\mathrm{sim}}\approx\left.\frac{\Delta T_{\mathrm{sim}}}{T_{\mathrm{g}}}\right|_{\mathrm{CuMn}}\,\left[\frac{\xi_{\mathrm{CuMn}}(t_{\mathrm{w}})}{\xi_{\mathrm{micro}}(t_{\mathrm{w}})}\right]^{1/(D_{\text{s}}/2-y)}\,,$ (73)
where the scaling law is based on Eq. (70). Consider two cases that differ in the cold temperature $T_{2}$: $T_{2}=0.5$ meets the chaos requirement of Eq. (73); $T_{2}=0.7$ does not [recall that the glass temperature is $T_{\mathrm{g}}=1.102(3)$].
The main result of Baity-Jesi et al. [31] is reproduced in Fig. 49. Rejuvenation is well visible in the central-bottom panel, as the peak of $S_{\mathrm{ZFC}}$ occurs at time $t^{\mathrm{eff}}_{\mathrm{w}}\ll(t_{\mathrm{w1}}+t_{\mathrm{w2}})$. This analysis is phenomenological and closely related to the experimental one, but the advantage of “numerical experiments” is that they give access to the full configurations and, hence, to microscopic correlation functions. To explain this phenomenon, we need three length scales: $\xi_{\mathrm{micro}}$, $\zeta$ and $\xi_{\mathrm{Zeeman}}$ (see Sec. I.3). Let us focus on Fig. 50. In a fixed-temperature protocol (native runs) $\xi_{\mathrm{Zeeman}}$ follows the behavior of $\xi_{\mathrm{micro}}$, recall Fig. 32. In contrast, in a jump run, the situation is more intricate. The size of the glassy domains, i.e., $\xi_{\mathrm{micro}}$, is frozen, but the system responds to the external field with a shorter $\xi_{\mathrm{Zeeman}}$ (rejuvenation). In particular, if the cold temperature $T_{2}$ satisfies Eq. (73), we have full chaos and $\xi_{\mathrm{Zeeman}}^{\mathrm{jump}}(T_{2},t_{\mathrm{w2}})\sim\xi_{\mathrm{Zeeman}}^{\mathrm{native}}(T_{2},t_{\mathrm{w2}}-t_{\mathrm{w1}})$; otherwise, we have an intermediate situation, as the $T_{2}=0.7$ case shows. The last ingredient is that $\xi_{\mathrm{Zeeman}}^{\mathrm{jump}}$ values can be matched by a simple factor to the values of $\zeta^{\mathrm{jump}}$. This implies that, in the presence of strong chaos, the defects are the ones responding to the magnetic field. This observation will be vital for the quantification of memory, as it allows us to perform numerical experiments in absence of the external field and simply calculate $\zeta$ to quantify $\xi_{\mathrm{Zeeman}}$ in temperature-varying protocols.
Physical intuition suggests that temperature chaos is key for understanding the origin of memory and rejuvenation. In Eq. (73), Baity-Jesi et al. [31] implicitly used the main results of Zhai et al. [301], Baity-Jesi et al. [30] to translate what a strong chaotic regime is from experiment to simulation. $T_{2}=0.5$ meets the requirement for strong chaos, while $T_{2}=0.7$ does not.
To quantitatively check this statement, the authors computed the chaotic parameter locally in a similar manner as explained in Sec. VI.1.2. However, the authors did not impose the strong criterion of $\xi_{\text{micro},T_{1}}(t_{\mathrm{w1}})=\xi_{\text{micro},T_{2}}(t_{\mathrm{w2}})$, but just took the configurations with the largest coherence length $\xi_{\mathrm{micro}}$ for each temperature, resulting in the largest chaotic signal ever reported in a numerical simulation. The probability distributions of Eq. (69) were computed with native configurations at $T_{1}=0.9$ and jump configurations at $T_{2}=0.5,0.7$. Fig. 51 shows that the $T_{2}=0.5$ curve exhibits strong chaos while the $T_{2}=0.7$ curve does not.
These results point to temperature chaos as a necessary condition for rejuvenation.
VI.2.3 Simulation and experiment: the memory coefficients
This section introduces a numerical observable for quantifying memory in temperature-varying protocols. It will be shown that, if the relevant length scales are known, a simple scaling law can describe all the different memory coefficients (numerical and experimental).
The equilibrium non-linear susceptibility is proportional to the integral of $r^{2}C_{4}$ for $0<r<\infty$ [63]. Thus, following Baity-Jesi et al. [36, 33], Zhai et al. [302], Paga et al. [232], one can generalize this equilibrium relation by computing these integrals using the non-equilibrium correlation function $C_{4}$ of Eq. (22). As shown in Fig. 52, there is a small but detectable difference in the behavior of $r^{2}C_{4}$ as $t_{\mathrm{w2}}$ varies in temperature cycles. Paga et al. [230] considered the curve with $t_{\mathrm{w2}}=0$ as the reference curve and proposed as numerical memory coefficient
$\mathcal{C}_{\mathrm{num}}=1-\frac{\int_{0}^{\infty}\mathrm{d}r\,r^{2}\,\Delta C_{4}(r,t_{\mathrm{w2}};p)}{\int_{0}^{\infty}\mathrm{d}r\,r^{2}\,C_{4}(r,0;p)}\,,$ (74)
where $\Delta C_{4}(r,t_{\mathrm{w2}};p)=C_{4}(r,t_{\mathrm{w2}};p)-C_{4}(r,0;p)$ and $p$ denotes the different protocols.
Figure 51: Comparison of probability distributions for the chaotic parameter defined in Eq. (69), recall Sec. VI.1.2, computed with the $T_{1}=0.9$ native protocol and the $T_{2}=0.5$ jump protocol (yellow curve) or the $T_{2}=0.7$ jump protocol (blue curve). Strong chaos correlates with the rejuvenation phenomenon. Figure taken from Baity-Jesi et al. [31].
Given the accuracy achieved for $\Delta C_{4}(r,t_{\mathrm{w2}},p)$, $\mathcal{C}_{\mathrm{num}}$ is sensitive to even tiny differences in the state of the system just before and just after the temperature cycle.
The natural choices for the scaling variables are
$x\!=\!\Big{[}\frac{\zeta(T_{2},t_{\mathrm{w2}})}{\xi_{\mathrm{micro}}^{\mathrm{native}}(T_{1},t_{\mathrm{w1}})}\Big{]}^{D-\theta/2}\!\!,\ y=\frac{T_{1}}{T_{\mathrm{g}}}\,\zeta(T_{2},t_{\mathrm{w2}})\,.$ (75)
Both $x$ and $y$ are approximately accessible to experiment through $N_{\text{s}}^{\text{jump}}$ and $N_{\text{s}}^{\text{native}}$. Paga et al. [230] proposed a simple scaling law,
$\mathcal{F}(x,y)=y\,[1+a_{1}x+a_{2}x^{2}]\,,$ (76)
which, with fitted $a_{1}$ and $a_{2}$ coefficients, describes the behavior of both experimental and numerical memory coefficients. The scaling representation, dashed line in Fig. 53 (top), shows that, away from the $\mathcal{C}\approx 1$ region, the relation $\mathcal{C}_{\chi^{\prime\prime}}\approx[\mathcal{C}_{\text{Zeeman}}]^{K}$ holds with $K\approx 3.9$. This shows that the different memory coefficients carry the same physical information. Although the scaling in Eq. (76) could be feared to be accurate only near ${\cal C}\approx 1$, the ansatz turns out to cover all reachable values of ${\cal C}$.
In these sections, a quantitative formulation has been developed for memory in rejuvenated glasses utilizing measured and calculated lengths (both correlation lengths and coherence lengths) and confirming that rejuvenation and temperature chaos are strongly related.
VI.3 A different kind of memory: the Mpemba effect
Figure 52: Top: the curves $r^{2}C_{4}(r,t_{\mathrm{w2}};p)$ vs $r$. Bottom: as in Top, but subtracting the native $C_{4}$ from the correlation function of the temperature-cycled system. As the time spent at the colder temperature $T_{2}$ increases, the system loses memory. The memory loss is evinced by the increasing signal in $r^{2}\Delta C_{4}$. Figure taken from Paga et al. [230].
Figure 53: Top: Experimental memory coefficients $\mathcal{C}_{\chi^{\prime\prime}}$ (circles) and $\mathcal{C}_{\mathrm{Zeeman}}$ (rhombus and squares) vs the scaling function in Eq. (76). Here, $x^{\prime}$ and $y^{\prime}$ are experimental proxies for the $x$ and $y$ scaling variables of Eq. (75). Empty circles refer to $t_{\mathrm{w1}}=1\,\mathrm{h}$ (full circles for $t_{\mathrm{w2}}=3\,\mathrm{h})$. The dotted line is a fit to a straight $\mathcal{C}_{\mathrm{Zeeman}}=1-\alpha\mathcal{F}^{\prime}$. The dashed line is $(1-\alpha\mathcal{F}^{\prime})^{3.9}$ and represents the experimental findings for $\mathcal{C}_{\chi^{\prime\prime}}$ surprisingly well. Bottom: As in Top, for coefficient $\mathcal{C}_{\mathrm{num}}$. See Tab. 4 for parameter details. Figure taken from Paga et al. [230]; data for the $\mathcal{C}_{\chi^{\prime\prime}}$ coefficient from Freedberg et al. [125].
$T$-drop $T_{1}$ $T_{2}$ $t_{\mathrm{w1}}$ $\xi_{\mathrm{micro}}(T_{1},t_{\mathrm{w1}})$
J1 1.0 0.625 $2^{19.5}$ 8.038(1)
J2 1.0 0.625 $2^{22}$ 10.085(15)
J3 1.0 0.625 $2^{24.625}$ 12.75(3)
J4 1.0 0.625 $2^{27.25}$ 16.04(3)
J5 1.0 0.625 $2^{28.625}$ 18.08(5)
J6 1.0 0.625 $2^{29.875}$ 20.20(8)
J7 1.0 0.700 $2^{29.875}$ 20.20(8)
J8 0.9 0.500 $2^{31.25}$ 16.63(5)
J9 0.9 0.700 $2^{31.25}$ 16.63(5)
Table 4: Parameters of the numerical simulations shown in Fig. 53. Table taken from Paga et al. [230].
Experimental and numerical spin glasses exhibit strong memory effects that are deeply connected with the aging state of the system (or equivalently, the characteristic length scales of the system) as has been explained above. This section follows Baity-Jesi et al. [35] to present a different kind of memory: the Mpemba effect.
Consider two beakers of water that are almost identical except for the fact that one is hotter than the other. If we use a thermal reservoir (for example, a freezer) at a temperature below freezing point, it can be observed under some circumstances that the initially hotter beaker freezes faster than the colder one. This is known as the Mpemba effect [216].
The interesting story of the discovery of this effect by Mpemba can be found in Mpemba and Osborne [216], although versions of this phenomenon have been reported since ancient history [19] and have been widely studied in systems like nanotube resonators [135], clathrate hydrates [4], granular fluids [187], and colloidal systems [185].
Figure 54: Time evolution of the energy of two spin-glass systems, one initially at a higher temperature $T=1.3$ and the other at a lower temperature $T=1.2$, both quenched to $T=0.7\approx 0.64T_{\mathrm{g}}$. The initially hotter system (i.e., having a higher energy $e$) colds faster than the initially colder (lower-energy) one in agreement with the original Mpemba experiment. Inset: Closeup of the crossing between energy-time curves. Error bars are equal to the thickness of the lines. Figure from Baity-Jesi et al. [35].
Baity-Jesi et al. [35] observed and explained this effect in a well-controlled system by taking advantage of numerical simulations in spin glasses. The authors simulated a 3D EA model in a lattice of system size $L=160$. As long as simulations are performed out of equilibrium, it is necessary to identify what faster cooling means, since temperature is only well defined in equilibrium. The natural candidate to take the place of the temperature seems to be the energy density $e(t)$ because it is the observable conjugate with (the inverse of the) temperature.
Two different off-equilibrium systems are simulated. In the first one, the temperature of the thermal reservoir is quenched from infinite temperature to $T_{1}=1.3$, which is above the critical temperature [$T_{\mathrm{g}}=1.102(3)$, see Baity-Jesi et al. [27]]. The system is left to evolve until it reaches the energy $e_{1}\approx-1.6428$ and this is the starting point of the first system in the numerical experiment. The second simulated system is prepared in a very similar way but the temperature of the thermal reservoir is now $T_{2}=1.2$ and the initial energy is lower, $e_{2}=-1.67174$.
At this point, the temperature of the thermal reservoir of both systems is quenched to $T=0.7\approx 0.64T_{\mathrm{g}}$, and the energies are recorded for several times. The results of this numerical experiment are shown in Fig. 54.
The key to understanding what is happening is the control parameter of the phenomenon: the coherence length. Although the hot-start system has a higher energy density in the initial setup, it is also older in the aging sense of the coherence length $\xi_{\text{micro},1}(t_{\mathrm{w}}=0)=12$. The cold-start system has a lower initial energy density but it is younger, with a coherence length $\xi_{\text{micro},2}(t_{\mathrm{w}}=0)=5$.
Two ingredients are responsible for the Mpemba effect in spin glasses. The first is the temperature dependence of the growth of the coherence length; for the system initially at $T=1.3$, $\xi_{\mathrm{micro}}$ grows faster than in the system at initial temperature $T=1.2$. The second one is that, while the coherence length $\xi_{\mathrm{micro}}$ is a slow variable that remains essentially constant when changing the temperature, the energy density is a fast variable and, as a first approximation, when a quick temperature change takes place, it changes instantaneously to the value of the energy density corresponding to its new thermal reservoir.
VII Issues for future investigations
We collect here a few unanswered questions whose answers, in our opinion, could lead to significant progress. We shall distinguish the more theoretical aspects (VII.1) from questions of an experimental nature (VII.2).
VII.1 What is yet to be nailed down theoretically
On the low-temperature phase.
The renormalization group (RG), which has served us so well to clarify the nature of phases and phase transitions in clean systems [291, 77, 14], needs to be extended to spin glasses. This will not be easy. Numerical simulations clearly hint that universality (perhaps the most important consequence of the RG framework) extends to disordered systems [39, 126, 128, 127], including classical spin glasses [145, 107]. For quantum spin glasses we even have direct experimental confirmation [177] of critical exponents obtained with numerical simulations [51].
Furthermore, it is very difficult to determine from experimental and numerical analysis the different universality classes, due to the large uncertainties in the values of the critical exponents. One of the exceptions to this situation was the computation of the critical exponents in the three-dimensional Ising Edwards-Anderson model by Baity-Jesi et al. [27]. As discussed in detail in this review, the Ising Edwards-Anderson model beautifully describes the non-equilibrium dynamics of CuMn samples, even if CuMn is a Heisenberg material. This discrepancy can be reconciled by noticing that small but unavoidable anisotropies in the spin-spin interactions lead to Ising universality [37]. However, not all is understood, since experimental measurements for critical exponents in spin glasses are not close to the Ising values. For example, the non-linear magnetization exponent $\gamma=6.33(11)$ was extracted from simulations of the three-dimensional Ising spin glass [27]. Probably the most accurate experimental determination was obtained from zero-field measurements on the Heisenberg spin glass 0.5 at.AgMn [190] and yielded a very different $\gamma=2.3\pm 0.2$. The authors are unaware of other experimental reports for values of $\gamma$ that come close to the value obtained from simulations. Similar differences between simulation and experiment are present for other critical exponents (e.g., $\beta,~{}\delta,~{}{\text{and}}~{}\nu$).
On the other hand, our favorite tool, the perturbative RG, does not lead to unambiguous predictions. Even seemingly simple questions still lack an answer. Think, for instance, of the stability of the spin-glass phase under an externally applied magnetic field (see Sec. III.2).
Some new ideas to carry out non-perturbative computations would clearly be welcome. These might include further development of real-space tools such as the Monte Carlo RG [241] or the $M$-expansion around the Bethe lattice [8, 17]. More speculatively, we could think of the metastate —the mathematically rigorous way of building the low-temperature phase [6, 222, 223, 224, 249], see Newman et al. [221] for a recent review. It is conceivable that the De Dominicis RG program at zero field  [98, 95] could blend with the metastate approach to build a rigorous theory of the spin-glass phase [249].
On numerical simulations.
As discussed in Sec. II.7, NP-completeness is making spin glasses increasingly popular in an engineering context. This has driven the development of specific hardware to solve spin-glass problems. One may easily speculate that the main difficulties hampering spin-glass numerical simulations are common to both the physics and the engineering contexts. It follows that eventual advances in simulation algorithms (or hardware) in physics may have far-reaching consequences.
Progress in this problem has been absent for the last two decades. Our best algorithm for sampling equilibrium configurations, parallel tempering, was introduced in the 1990s [129, 152, 270]. The largest system equilibrated deep in the spin-glass phase in $D=3$ has been an $L=32$ cubic lattice for 15 years [12]. We have some evidence hinting to a physical origin for these difficulties, namely temperature chaos [111, 58].
Beside new hardware developments (see Sec. II.7), new simulation algorithms are needed. Temperature chaos suggests that these should work at a single temperature. Alternatively, ingenious ways of circumventing chaotic phenomena should be devised. Quantum annealing could be such an approach, but it may be hampered by chaotic phenomena related to the transverse field [180].
Experiments with potential for strong impact on theory.
As explained in Sec. IV.3, important experiments in the 1990s and the 2000s convinced the community about the far-reaching nature of the generalized fluctuation-dissipation relations. In the last ten years, single-crystal samples supporting large spin-glass coherence lengths $\xi_{\mathrm{micro}}$ have been obtained. A more detailed experimental investigation of the fluctuation-dissipation relations at larger $\xi_{\mathrm{micro}}$ would be clearly welcome.
Experiments on an ensemble of small spin-glass grains, where $\xi_{\mathrm{micro}}$ would be of the order of the grains’ size, open up another avenue. Under these conditions, large sample-to-sample fluctuations may unveil exotic phenomena such as multifractality [32]. Reliably measuring single-grain magnetizations would be a (not minor) technical challenge but even the collective magnetization of a (small) set of grains may be useful to investigate finite-size effects experimentally. This will bring us closer to the broader field of structural glasses, where non-linear responses unveil length scales akin to $\xi_{\mathrm{micro}}$ [1, 7].
VII.2 Experimental opportunities and challenges
The remaining issues that experiments can address arise from our current understanding of spin-glass dynamics. As is so often the case with spin glasses, there will be surprises that will dictate other opportunities for investigation. And, of course, the interaction with simulations and theoretical advances will suggest even other lines of inquiry. For now, we briefly describe a few of the more obvious directions for experimental research.
The onset of temperature chaos.
Temperature chaos is thought to be the origin of rejuvenation (see Sec. VI.2), but the question of its onset remains. Is it a sharp or broad transition as a function of temperature change? Is there a mixture of reversible and chaotic states in the transition region? The jury is still out on these fundamental issues. The experiments of Guchhait and Orbach [139] exhibit a finite temperature window of about 100 mK between complete reversibility and complete chaos, while those of Zhai et al. [301] find a very narrow temperature range of 10 mK for the onset of temperature chaos. Further, the latter find evidence for enhanced chaos as the temperature change increases. The two experiments are on different systems, but the question remains: what happens in the transition region?
Probing thin-film dynamics for times greater than $t_{\text{co}}$.
As pointed out in Sec. V.1, experiments on thin films were interpreted in terms of two time regimes. The first was the growth of the correlation length from nucleation to the film thickness $\mathcal{L}$ at a time defined as $t_{\text{co}}$, with a rate of growth appropriate to 3D. The second was an apparent instantaneous value reached in the parallel direction, appropriate to 2D. The simulations of Sec. V.1 however, found four time regimes for the growth of the correlation length. In particular, after the correlation length reached the vicinity of $\mathcal{L}$, the correlation length continued to grow more rapidly than in 3D, finally reaching an equilibrium value appropriate to 2D. It would be of interest to measure the magnitude of the correlation length experimentally well after $t_{\text{co}}$. This will require much longer time measurements than have been taken heretofore.
The nature of rejuvenation.
Rejuvenation is nearly always attributed to temperature chaos. Is this true? An alternate view has been proposed by Berthier and Bouchaud [53, 54], who “demonstrated in a numerical simulation of the 4D Ising spin glass that a temperature change $\Delta T/T_{\mathrm{g}}$ induces strong rejuvenation effects, whereas the direct observation of the overlap between configurations reveals no sign of chaos on the dynamically relevant length scales.” Making reference to the claim of Jönsson et al. [163] that temperature chaos is the origin of rejuvenation, they state that this analysis “cannot be viewed as definitive evidence for temperature chaos.” Note that in Sec. 41 we point out that Jönsson et al. [163] obtained a value $1/(D_{\text{s}}/2-y)\approx 2.6$ whereas Berthier and Bouchaud [53] use $\/(D_{\text{s}}/2-y)\approx 1$, the value found in most simulations and by Zhai et al. [301]. This issue may be illuminated by the experiments suggested above.
Spin glasses in a field: crossover or condensation?
The vexing question of whether there is a spin-glass transition in the presence of a magnetic field, and the allied question of whether the de Almeida-Thouless line [93], predicting a magnetic-field-dependent transition temperature $T_{\mathrm{g}}(H)$, exists has bedeviled theorists and experimenters alike. This issue was discussed in Sec. III.2, but we briefly restate it.
The mean-field theory predicts a phase transition to an ordered spin-glass phase for $T<T_{\mathrm{g}}(H)$. The droplet picture, however, finds that “a magnetic field destroys the spin-glass phase so that there is no de Almeida-Thouless transition” [112]. The long-distance dependence of the correlation function between spins located at position $i$ and $j$ is $C_{ij}\sim q_{\text{EA}}^{2}T/(Y|i-j|^{\theta})$, with $\theta\simeq 0.2$ for $D=3$, and $Y$ a generalization of the twist modulus or inter-facial tension. This very slow fall off may explain why the spin-glass system appears critical for $T<T_{\mathrm{g}}$.
Thus, there are two very different pictures for the behavior of a spin glass in the presence of a magnetic field. The numerical picture for $D=3$ is muddled (Sec. V.5), while a transition has been found in higher dimensions (Sec. III.2). Experimentally, there is evidence for a de Almeida Thouless line [173] consistent with the mean-field picture. The fundamental question is whether this is indicative of a phase transition to a spin-glass state in the presence of a magnetic field, or simply a crossover to the droplet-style phase. A future experimental investigation of ac susceptibility compared to dc magnetization measurements may shed light on this fundamental issue.
Spin-glass magnetization in high magnetic fields.
With the development of high magnetic field facilities such as the High Magnetic Field Laboratory at Florida State University, it is possible to fully saturate a dilute spin glass that might give insight to the interactions between spins. There has been a preliminary unpublished effort on CuMn to explore such. In the simplest case, deviations from a Brillouin function could be modeled to determine the interactions. In the preliminary study, it would appear such an analysis could provide information on at least nearest-neighbor interactions. This work is in its infancy but the initial results look promising.
Dynamics of spin glasses using aging.
The ac susceptibility has been seen to be a powerful tool to investigate aging, rejuvenation and memory. It is not often discussed that these intrinsic spin-glass features disappear with increasing the frequency of ac susceptibility measurements. At least in the case of the recent work on CuMn it was found that aging and therefore rejuvenation and memory were unmeasurable at frequencies on the order of 10 Hz. Considering this in terms of an Arrhenius law, it implies there is a minimum energy barrier relevant to the spin-glass state. What has not been investigated is the frequency dependence of aging as a function of temperature. This would require measurements of both the real and imaginary parts of the susceptibility. For example, if one interprets the disappearance of aging as a measure of a minimum energy barrier then at the disappearance of aging the imaginary part would go to zero since, without transitioning over energy barriers, there would be no dissipation.
An opposite perspective would take note of the absence of rejuvenation and memory for measurement frequencies greater than 10 Hz; that is, for response times faster than $10^{-1}$ s. Because of ultrametric symmetry, the response time of the vast majority of states is set by the highest barrier created in a waiting time $t_{\mathrm{w}}$. “Asking” the spin glass to respond more rapidly than that time (i.e., shorter times) would involve states with lower barrier heights, but their number is exponentially smaller than those at the largest barrier height. Concomitantly, the response at a higher frequency would be exponentially smaller.
To test this perspective, one needs make measurements at different $t_{\mathrm{w}}$. For longer $t_{\mathrm{w}}$, the response time would be longer, so that one would have to go to lower frequencies to observe rejuvenation and memory.
VIII Summary and Conclusions
The authors have attempted to inform the reader of this review of the early history of spin glasses, the subsequent quantitative developments, both theoretical and experimental, coalescing into our current understanding of their many facets. We feel it is important to illustrate the dynamics of this “useless” material because of its remarkable general importance across a wide spectrum of scientific research. Indeed, the 2024 physics Nobel prize went in part to J.J. Hopfield for his application of the structure of spin-glass dynamics to neural networks.
We have led the reader through the very early experimental manifestations of spin glasses, well before their general nature was understood. The crucial experiments were presented that defined the nature of the spin glass. This was followed by a description of the various models developed to elucidate the nature of the spin-glass state. We did not include many important ancillary developments associated with spin-glass dynamics because of space limitations, but rather attempted to focus on a perhaps narrow perspective, but one that we believe provides a rather complete picture within its boundaries.
We hope that the reader will be encouraged by our presentation to join with us in this pursuit of Anderson’s cornucopia of opportunities.
Acknowledgments
This work was partially supported by Ministerio de Ciencia, Innovación y Universidades (MICIU, Spain), Agencia Estatal de Investigación (AEI, Spain), and European Regional Development Fund (ERDF, A way of making Europe) through Grants no. PID2020-112936GB-I00 and PID2022-136374NB-C21. EM acknowledges the hospitality of Mark Bowick at UCSB (USA), where part of this work was done, and the support by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). His research was supported by the 2021 FIS (Fondo Italiano per la Scienza) funding scheme (FIS783 - SMaC - Statistical Mechanics and Complexity) and the 2022 PRIN funding scheme (2022LMHTET - Complexity, disorder and fluctuations: spin glass physics and beyond) from Italian MUR (Ministry of University and Research). EDD and RLO were supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Science and Engineering Division, under Award No. DE-SC0013599. The single CuMn crystals were grown by Dr. D.L. Schlagel at Ames Laboratory, which is operated by the U.S. DOE by Iowa State University under Contract No. DE-AC02-07CH11358. IGAP and FRT acknowledge the support of ICSC - Italian Research Center on High-Performance Computing, Big Data, and Quantum Computing, funded by the European Union - NextGenerationEU.